### v1

# Gated Rotary-Enhanced Linear Attention for Long-term Sequential Recommendation

Abstract.
In Sequential Recommendation Systems (SRSs), Transformer models show remarkable performance but face computation cost challenges when modeling long-term user behavior sequences due to the quadratic complexity of the dot-product attention mechanism. By approximating the dot-product attention, linear attention provides an efficient option with linear complexity. However, existing linear attention methods face two limitations: 1) they often use learnable position encodings, which incur extra computational costs in long-term sequence scenarios, and 2) they may not consider the user’s fine-grained local preferences and confuse these with the actual change of long-term interests. To remedy these drawbacks, we propose a long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention (RecGRELA). Specifically, we first propose a Rotary-Enhanced Linear Attention (RELA) module to model long-range dependency within the user’s historical information using rotary position encodings. We then introduce a local short operation to incorporate local preferences and demonstrate the theoretical insight. We further introduce a SiLU-based Gated mechanism for RELA (GRELA) to help the model determine whether a user’s behavior indicates local interest or a genuine shift in long-term preferences. Experimental results on four public datasets demonstrate that our RecGRELA achieves state-of-the-art performance compared to existing SRSs while maintaining low memory overhead.
Sequential Recommendation, Linear Attention, Rotary Position Encoding, Gated Mechanism
ccs: Information systems Recommender systems
1. Introduction
Figure 1. Illustration of differences of SASRec, LinRec, and our RecGRELA. $\phi(\cdot)$ denotes the kernel function. (a) SASRec utilizes Softmax attention, where similarities between all query-key pairs are computed, leading to quadratic complexity for sequence length. (b) LinRec implements linear attention by transforming Q and K with $\phi(\cdot)$. This reordering of computations achieves linear complexity, though it may diminish the model’s expressive power. (c) Our RecGRELA enhances linear attention by incorporating Rotary Positional Encoding (RoPE) for Q and K and adding a shortcut connection, aiming to achieve strong expressive capability while maintaining linear computational complexity and removing the need for explicit Learned Position Encodings (LPE).
Sequential Recommender Systems (SRSs) (Wang et al., 2019; Fang et al., 2020; Pan et al., 2024; Niu et al., 2025; Zhang et al., 2025; Yang et al., 2023; Zhang et al., 2021) have become a fundamental technology in personalized services, significantly transforming how online platforms predict user behavior. By analyzing sequences of user interactions across e-commerce, streaming services, and social media, SRSs capture both short-term interests and long-term preference trends. However, modeling long-term dependencies is challenging, as efficiently handling long sequences demands a balance between computational cost and model performance, while also capturing the dynamic and complex patterns in user interests. To address the challenge of modeling user behavior sequences, researchers explore various methodologies. Traditional approaches, such as those based on Markov Chains (He et al., 2017; Rendle et al., 2010) or Recurrent Neural Networks (RNNs) (Chen et al., 2018; Cui et al., 2018; Donkers et al., 2017; Xu et al., 2019), lay the groundwork but also face limitations in capturing long-range dependencies or suffered from computational inefficiencies during training and inference, particularly with increasing sequence lengths. Consequently, the field moves towards more advanced models that can handle complex sequential dependencies. This also presents several new challenges for the field.
Specifically, the first challenge is Efficiency. Transformer models (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019) emerge as a compelling replacement for earlier approaches like RNNs, leveraging their ability to process sequential data and capture both local and global preferences of users. However, the dot-product attention (as show in Fig. 1(a)) in Transformers exhibits quadratic complexity relative to sequence length, resulting in high costs for long sequences. To address this issue, linear attention (Katharopoulos et al., 2020; Liu et al., 2023) has been proposed as an efficient alternative, providing linear complexity (as show in Fig. 1(b)). However, compared to dot-product attention, current linear attention approaches may suffer from performance drops.
In addition to the efficiency of attention mechanisms, effectively capturing the positional information in user interaction sequences is another crucial challenge that affects both model performance and operational efficiency. Position encoding is widely adopted to capture the sequential order of items users interact with due to its ability to enhance the positional awareness of contextual vectors. As illustrated in Fig. 1(a) and (b), the SASRec (Kang and McAuley, 2018) and LinRec (Liu et al., 2023) both rely on learnable position encoding. The learned positional encoding requires more parameters than fixed positional encoding schemes since each position needs its own learned vector, and they do not have any inherent structure that captures the notion of relative distances between positions. To handle these drawbacks, some works (Su et al., 2024; Tian et al., 2024) employ an improved Rotary Position Encoding (RoPE), which injects relative positional information in a more efficient manner. However, they still use the dot-product attention mechanism, which limits its scalability for long sequences. Furthermore, the RoPE exhibits inefficiency in multi-head attention scenarios, restricting the model’s overall expressive capability.
Linear attention, while effective for grasping overarching long-term dependencies, local preferences are equally vital for the long-term sequential recommendation, driven by users’ immediate preferences. For instance, a user might intensely engage with a specific series of related items. They could watch multiple episodes of The Avengers and then immediately consume supplementary content detailing Marvel’s production process. Subsequently, they might delve into documentaries that explore the creators behind such cinematic universes. However, linear attention’s tendency to smooth features can obscure important local information, while RoPE’s distance-based decay (Su et al., 2024) weakens attention to nearby items, particularly in long sequences. Both these characteristics hinder the precise modeling of local patterns. Even when linear attention successfully captures both long-term interests and short-term local preferences, a significant challenge persists: when a user’s preferences really change, it’s hard for the model to identify them. For instance, consider a user like historical documentaries over a weekend, binge-watches several interconnected Marvel movies like The Avengers series, and then explores some behind-the-scenes content about their production. A model that struggles to differentiate signals might misinterpret this intense but potentially short-lived burst of activity as a long-term shift in preference. Consequently, it might aggressively recommend a deluge of Marvel-related content for weeks – from other superhero films they’ve already seen, to comic book adaptations, or even tangentially related action movies. This happens because the model overweights fleeting interests. The user might simply want a weekend of superhero fun and is now looking for a historical drama, getting frustrated by the repetitive and misaligned suggestions.
To solve the aforementioned challenges, we propose RecGRELA, a novel long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention block (as shown in Fig. 1(c)) to effectively capture long-range dependency. Specifically, we design a Rotary Enhanced Linear Attention (RELA) module. RELA builds upon linear attention, efficiently solving the computation limitation with a linear complexity of $\mathcal{O}(N)$, It also enhances its positional awareness by incorporating Rotary Position Encoding (RoPE). RoPE enables the model to capture relative positional relationships between items in a parameter-efficient manner in multi-head scenarios, overcoming the drawback of learnable positional encodings that introduce additional parameters. Concurrently, we design a Local Shortcut operation with causal 1D convolution in RELA to integrate local preferences and present theoretical insights based on the matrix rank analysis. We further introduce a SiLU-based gated mechanism to identify whether user behavior is a temporary burst of interest or a true shift of long-term interest via adjusting the output of RELA. With this gated mechanism, we remove the linear projection on the value matrix of the attention module to reduce the computational cost.
The contributions of this work are summarized as follows:
  • We propose a Rotary Enhanced Linear Attention (RELA) module that integrates Rotary Position Encoding (RoPE) into the linear attention mechanism to address efficiency issues in long-sequence modeling. By replacing dot-product attention with linear attention, our approach reduces computational costs while maintaining the ability to capture inter-item positional relationships.
  • We further develop a Local Shortcut operation to inject dense local interaction into the linear attention framework. This mechanism combines causal 1D convolution, explicitly modeling local preferences of user behaviors without sacrificing computational efficiency. We introduce a SiLU-based gated linear attention mechanism to distinguish true shifts in user preferences from transient local interests.
  • Based on these designs, we further construct the RecGRELA, a novel long-term sequential recommendation paradigm that achieves competitive performance compared with state-of-the-art SRSs on four real-world datasets, keeping low memory overhead. Our source code is available at Github111 https://github.com/TBI805/RecGRELA..
Table 1. Notations and their descriptions for Section 2.
Notations Description
$\mathcal{U}$ user set with each user $u\in\mathcal{I}$
$\mathcal{V}$ user set with each user $v\in\mathcal{V}$
$\mathcal{S}_{u}$ interaction sequence for user $u$
$N$ sequence length
$D,d$ embedding dimension and head dimension of training sequence
$\boldsymbol{X}$ input sequence
$\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}$ projection matrix for $\boldsymbol{X}$
$\boldsymbol{Q}_{m},\boldsymbol{K}_{m},\boldsymbol{V}_{m}$ query, key and value matrix on position $m$
$\boldsymbol{O}_{m}$ attention result on position $m$
$\mathcal{O}\left(\cdot\right)$ computation complexity
$\phi\left(\cdot\right)$ kernel function
$\mathbf{R}$ the rotation matrix
$\mathcal{C}\left(\cdot\right)$, $\mathcal{C}^{-1}\left(\cdot\right)$ operator for converting between real and complex vector representations
$\mathcal{R}\left(\cdot\right)$ RoPE operation
$\mathrm{Re}[\cdot]$ real part of the complex vector
2. PRELIMINARY
2.1. Problem Definition
[…]
2.2. Dot-Product Attention and Linear Attention
Let $\boldsymbol{X}\in\mathbb{R}^{N\times D}$ denote an input sequence. Formally, the input sequence $\boldsymbol{X}$ is projected by three projection matrices $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}\in\mathbb{R}^{D\times d}$, to corresponding representations $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}_{Q}$, $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}_{K}$, and $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}_{V}$. The dot-product attention (Vaswani et al., 2017) can be written as:
(1) $\quad{\boldsymbol{O}}_{m}=\sum_{n=1}^{N}\frac{\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}{\sum_{n=1}^{N}\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}\boldsymbol{V}_{n},$
where $\boldsymbol{Q}_{m},\boldsymbol{K}_{n},\boldsymbol{V}_{n}\in\mathbb{R}^{1\times d}$, and $d$ is the head dimension. In this case, calculating attention weights in dot-product attention leads to the computation complexity of $\mathcal{O}(N^{2})$ with respect to the input sequence length $N$. For autoregressive modeling, dot-product attention can be implemented by incorporating a masking mechanism, which restricts the influence on the $m$-th position to a position $n$ satisfying $n\leq m$.
Linear attention (Katharopoulos et al., 2020) is proposed as an effective alternative that reduces the computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ compared with dot-product attention. Specifically, linear attention modifies the dot-product attention by replacing the non-linear Softmax operation with linear normalization, while incorporating an additional kernel function $\phi$ in query matrix $\boldsymbol{Q}$ and key matrix $\boldsymbol{K}$:
(2) $\boldsymbol{O}_{m}=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}{\sum_{n=1}^{N}\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}\boldsymbol{V}_{n}.$
By utilizing the associative property of matrix multiplication, the computation is reordered from $(\boldsymbol{Q}\boldsymbol{K}^{\top})\boldsymbol{V}$ to $\boldsymbol{Q}(\boldsymbol{K}^{\top}\boldsymbol{V})$:
(3) $\boldsymbol{O}_{m}=\frac{\phi\left(\boldsymbol{Q}_{m}\right)\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
where the computational complexity is decreased to $\mathcal{O}(N)$. In (Katharopoulos et al., 2020), the kernel function employs the Exponential Linear Unit (ELU) activation (Clevert, 2015), which prevents gradient vanishing when processing negative values of $\boldsymbol{X}$. In our work, we follow this setting for a fair comparison.
2.3. Position Encoding
Absolute Position Encoding (APE) (Vaswani et al., 2017) incorporates fixed values into the embeddings of sequence elements, establishing a consistent reference frame across all tokens. For a token at the position $t$ within a sequence of length $N$, this encoding is computed through a deterministic function utilizing paired sine and cosine operations across different frequencies. The formulation is denoted as follows:
(4) $PE\left(p,2t\right)$ $=\sin\left(p/10000^{2t/\left(D/2\right)}\right),$
$PE\left(p,2t+1\right)$ $=\cos\left(p/10000^{2t/\left(D/2\right)}\right),$
where p is the position of the item, t is the dimension of p.
Learnable Position Encoding (LPE) (Wang et al., 2021; Kang and McAuley, 2018) adopts a distinct methodology by modeling positional information as learnable parameters that are optimized during the training process. Specifically, for an item embedding x at position $m$, the transformation is performed as follows:
(5) $\textbf{{x}}_{m}=\textbf{{x}}+p_{m},$
where $p_{i}$ denotes the APE or LPE.
Rotary Position Encoding (RoPE) (Su et al., 2024) captures relative positional information (Shaw et al., 2018) through absolute position encoding methods. To achieve this, RoPE encodes the relative positional information via a rotation matrix $\mathbf{R}\in\mathbb{C}^{N\times(d/2)}$:
(6) $\mathbf{R}\left(m,t\right)=e^{i\theta_{t}m},$
[…]
(7) $\bar{\boldsymbol{Q}}^{\prime}_{m}$ $=\boldsymbol{\bar{Q}}_{m}\odot\mathbf{R}\left(m,t\right),$
$\bar{\boldsymbol{K}}^{\prime}_{n}$ $=\boldsymbol{\bar{K}}_{n}\odot\mathbf{R}\left(n,t\right),$
$\boldsymbol{A}_{(m,n)}^{\prime}$ $=\mathrm{Re}[\bar{\boldsymbol{Q}}^{\prime}_{m}\bar{\boldsymbol{K}}^{\prime}_{n}]=\mathrm{Re}[\boldsymbol{\bar{Q}}_{m}\boldsymbol{\bar{K}}_{n}^{*}e^{i(m-n)\theta_{t}}],$
where $\mathrm{Re}[\cdot]$ represents real part of the complex vector and * means complex conjugates, and $\odot$ denotes Hadamard product. Note that the attention matrix $\boldsymbol{A}_{(m,n)}$ with RoPE implies relative position in rotation form $e^{i(m-n)\theta_{t}}$ for $(D/2)$ frequencies. We can define the RoPE operation $\mathcal{R}\left(\cdot\right)$ to an input real vector $\boldsymbol{X}$ as follows:
(8) $\mathcal{R}\left(\boldsymbol{X}\right)=\mathcal{C}^{-1}\left(\mathcal{C}\left(\boldsymbol{X}\right)\odot\mathbf{R}\right),$
where $\mathcal{C}\left(\cdot\right)$ converts a real-valued tensor into a complex-valued tensor, and $\mathcal{C}^{-1}\left(\cdot\right)$ converts a complex-valued tensor into a real-valued tensor.
Table 2. Notations and their descriptions for Section 3.
Notations Description
$\boldsymbol{E}$ learnable embedding matrix
$\boldsymbol{H}_{o}$, $\boldsymbol{H}$ initial and final embedding matrix
$\boldsymbol{H}_{g}$ input of the the GRELA block
$\boldsymbol{H}_{L}$ final output for prediction
$\boldsymbol{O}_{f}$ output of the RELA
$\boldsymbol{S}$ output of the causal convolution
$\operatorname{CSC}(\cdot)$ causal convolution operation
$\boldsymbol{M}_{\operatorname{CSC}}$ sparse matrix associated with causal convolution and non-linear activation function
$\boldsymbol{M}_{eq}$ equivalent full attention map
$\boldsymbol{W}_{1}$, …, $\boldsymbol{W}_{5}$ weight matrix for linear projection
$\boldsymbol{b}_{1},...,\boldsymbol{b}_{5}$ bias for linear projection
$\boldsymbol{O}_{i}$ attention result on position $i$
$\hat{\mathcal{Y}}$ predicted recommendation result
$\mathcal{L}_{CE}$ cross-entropy loss
3. METHODOLOGY
In this section, we introduce our proposed framework, RecGRELA. The framework overview is shown in Figure 2.
Figure 2. The overview of our proposed architecture. The Overall Architecture (left) processes an input sequence through an Embedding Layer, which is then fed into a stack of $L$ identical encoder layers. Each encoder layer consists of Layer Normalization, a GRELA Block, another Layer Normalization, and an MLP, finally leading to a Prediction Layer. The central components are the GRELA Block (middle-bottom), which utilizes RoPE (Rotary Position Encoding, middle-top), and the RELA module (right). Within the RELA module, query ($Q$) and key ($K$) are processed by ELU activation and the RoPE. Attention scores are computed via matrix multiplication of the RoPE-enhanced $Q$ and $K$ (after $K$ is scaled), followed by another scaling and a concatenation step. A Causal Conv1D and SiLU activation processes the value ($V$) projection to capture local context.
The presentation is structured as follows: initially, we introduce the embedding layer; secondly, we elaborate on the core components of the GRELA Block; subsequently, we present the details of the Multi-Layer Perceptron, and finally, we describe the prediction layer and training strategies. The key notations of this section are listed in Table 2.
3.1. Embedding Layer
Following (Kang and McAuley, 2018), we utilize an embedding layer to map the training sequence $\mathcal{S}_{u}$ into a high-dimensional space. The embedding layer uses a learnable embedding matrix $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $\mathcal{V}$ denotes the item set and $D$ represents the embedding dimension. Given an input interaction sequence $\mathcal{S}_{u}$ of length $N$, it can be projected into the initial embedding matrix $\boldsymbol{H}_{o}\in\mathbb{R}^{N\times D}$ as follows:
(9) $\boldsymbol{H}_{o}=\boldsymbol{E}\mathcal{S}_{u}.$
Then we apply the Dropout (Srivastava et al., 2014) and layer normalization (Lei Ba et al., 2016) operations to obtain the final embedding $\boldsymbol{H}\in\mathbb{R}^{N\times D}$:
(10) $\boldsymbol{H}=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\boldsymbol{H}_{o}\right)\right).$
Attention-based sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020; Liu et al., 2023) typically inject a learnable position encoding into the final embedding, while our method captures position awareness with rotary position encoding by multiplying the rotation matrix with outputs of the kernel function in linear attention. This design eliminates the need to incorporate positional information into the embedding layers.
3.2. GRELA Block
In this part, we present the design of our proposed GRELA Block, which includes the Rotary-Enhanced Linear Attention module and the Local Shortcut operation.
3.2.1. Rotary-Enhanced Linear Attention
Overall Design. Previous work (Liu et al., 2023) provides $\mathcal{O}(N)$ computational complexity of attention mechanism to alleviate the issue of quadratic complexity in dot-product attention for sequential recommendation. However, it ignores the high cost of learnable position encoding in long-sequence modeling. Meanwhile, some works (Tian et al., 2024; Lopez-Avila et al., 2024) notice these issues and propose improved RoPE to address them, but unfortunately, they fail to address the problem of dot-product attention with $\mathcal{O}(N^{2})$ complexity. To bridge this gap, we propose a novel Rotary-Enhanced Linear Attention (RELA) mechanism, aiming to simultaneously resolve the issues of quadratic complexity and ineffective position encoding in long-sequence modeling. The motivation stems from the observation that other position encoding methods operate directly on the input sequence or the attention matrix. However, linear attention does not compute the attention matrix, which means other methods can only modify the input sequence. In contrast, RoPE achieves relative position encoding by leveraging absolute position encoding, by applying transformations only to the queries $\boldsymbol{Q}$ and keys $\boldsymbol{K}$. This design makes RoPE suitable for integration with linear attention. In a general situation, following Eq.(7), we revise Eq. (3) as follows:
(11) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
where $\mathcal{R}\left(\cdot\right)$ denotes injecting the RoPE into queries or keys. Following (Katharopoulos et al., 2020), we set $\phi(x)=\operatorname{ELU}(x)+1$. This kernel function ensures that the denominator remains positive to avoid the risk of division by zero, while the summation in the numerator also contains no negative terms.
Inspired by techniques for stabilizing gradients in dot-product attention (Vaswani et al., 2017), we further incorporate scaling factors. Specifically, the keys $\boldsymbol{K}$ and values $\boldsymbol{V}$ are scaled by $\sqrt{N}$ (where $N$ denotes the sequence length). This emulates the scaling mechanism used in conventional dot-product attention and helps regulate the dynamic range of the resulting key-value memory. Furthermore, the normalization term in the denominator, which is critical for linear attention, is constructed using the dot product between the query $\phi(\boldsymbol{Q}_{m})$ and the mean of keys $\phi(\boldsymbol{K}_{n})$. A small constant $\epsilon$ is added to this denominator for numerical stability, preventing division by zero and ensuring well-behaved gradients. This approach enables more fine-grained control over the attention distribution. The final RELA is formulated as follows:
(12) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)/\sqrt{N}\right)\left(\boldsymbol{V}_{n}/\sqrt{N}\right)\right)}{\phi(\boldsymbol{Q}_{m})\left(\frac{1}{N}\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}+\epsilon\right)}.$
In-Depth Analysis for Rotaty Position Encoding. To intuitively understand why Eq.(7) embodies the concept of Rotation, we introduce RoPE with the head dimension $d=2$. When the dimension $d=2$, we take $\boldsymbol{Q}_{m}$ as an example for a detailed derivation. We first express it in its complex form:
(13) $\boldsymbol{Q}_{m}=\left[\boldsymbol{Q}_{m}^{(1)},\boldsymbol{Q}_{m}^{(2)}\right]=\left[\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right],$
where $\boldsymbol{Q}_{m}^{(1)}$ and $\boldsymbol{Q}_{m}^{(2)}$ indicate the components of $\boldsymbol{Q}_{m}$ expressed in the 2D coordinates. Based on Euler’s formula, $e^{im\theta}$ can be rewritten as follows:
(14) $e^{im\theta}=\cos m\theta+i\sin m\theta.$
Then, we have:
(15) $\boldsymbol{Q}_{m}e^{im\theta}$ $=\left(\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right)(\cos(m\theta)+i\sin(m\theta))$
$=\left(\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)\right.$ $\left.-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta)\right)+i\left(\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)\right).$
Then, the result is re-expressed in the form of a real-valued vector:
(16) $\boldsymbol{Q}_{m}e^{im\theta}$ $=[\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta),\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)]$
$=\left(\begin{array}[]{cc}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{array}\right)\binom{\boldsymbol{Q}_{m}^{(1)}}{\boldsymbol{Q}_{m}^{(2)}}.$
Evidently, the matrix $\left(\begin{array}[]{cc}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{array}\right)$ describes a rotational transformation applied to the embedded sequence.
While the core rotational mechanism of RoPE effectively injects relative positional information, a critical design choice is the dimension upon which these rotations are applied. Conventionally, in multi-head attention architectures, RoPE is applied independently to each head’s query and key vectors, meaning the rotations operate on the head dimension $d$ (Su et al., 2024; Tian et al., 2024). This allows each head to potentially learn or leverage positional cues in its specialized subspace. Our choice deviates from this standard practice. We apply RoPE directly to the full embedding dimension $D$ of the queries and keys. The motivation is that imbuing the entire token representation with a unified positional signal at the embedding dimension $D$ level can lead to a more holistic and consistent understanding of sequence order throughout the model. Instead of each head independently re-interpreting or establishing positional context, all downstream computations operate on an input that is already position-aware across its full breadth.
3.2.2. Local Shortcut and Gated Mechanism
Attention mechanisms excel at long-range modeling, yet local behavior modeling is equally vital for long-term sequential recommendation systems, as user interactions often display strong local preferences (e.g., session-based behaviors) crucial for next-item prediction. However, linear attention variants falter here. Unlike Softmax attention’s sharp focus, linear attention’s smooth distribution tends to average features, obscuring crucial local cues (Han et al., 2023). Moreover, RoPE suffers from long-term decay (Su et al., 2024), meaning the attention score decreases as the relative distance increases. Because of linear attention’s smoothing and RoPE’s distance decay, precisely modeling crucial local patterns for long-term sequential recommendation becomes difficult. To address these issues, we introduce a Local Shortcut to mitigate the long-term decay of RELA. Specifically, we use a causal 1D convolution (Fu et al., 2023) followed by a SiLU activation (Ramachandran et al., 2017) to capture local preferences in the user interaction sequence:
(17) $\boldsymbol{S}=\operatorname{SiLU}\left(\text{CausalConv1d}\left(\boldsymbol{V}\right)\right),$
where $\boldsymbol{V}$ denotes the input sequence, and $\boldsymbol{S}$ represents the output of the temporal convolution. Then $\boldsymbol{S}$ is connected with the output $\boldsymbol{O}$ of linear attention as a shortcut to form the final output $\boldsymbol{O}_{f}$ of RELA:
(18) $\boldsymbol{O}_{f}=\boldsymbol{O}+\boldsymbol{S}.$
With the local shortcut, each query only focuses on several adjacent items instead of the whole sequence. This locality ensures that even if the linear attention values corresponding to two queries are the same, we can still get different outputs from different local preferences.
The theoretical motivation behind such a design concerns the upper bound on the rank of the attention matrix in linear attention (Han et al., 2023). The rank of the attention matrix is bounded by the length of the sequence $N$ and the feature dimension $d$ for each head:
(19) $\operatorname{rank}(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})^{\top})$ $\leq\min\{\operatorname{rank}(\phi(\boldsymbol{Q})),\operatorname{rank}(\phi(\boldsymbol{K}))\}$
$\leq\min\{N,d\},$
where $d$ is often smaller than $N$ in long-term sequence scenarios. Under these conditions, the upper bound on the rank of the attention matrix is constrained to remain at a low level. The rank represents the number of linearly independent vectors, i.e., the dimensionality, contained in the matrix. When an $N$-dimensional attention matrix has a rank of $d$, it effectively operates as a $d$-dimensional matrix in terms of its representational capacity, which fundamentally limits the expressive power of the attention mechanism. Referring back to Eq. 18, the operation on the attention matrix can be reformulated as:
(20) $\boldsymbol{O}_{f}$ $=\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\operatorname{CSC}(\boldsymbol{V})$
$=(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\boldsymbol{M}_{\operatorname{CSC}})\boldsymbol{V}=\boldsymbol{M}_{eq}\boldsymbol{V},$
where we omit the normalization term for simplicity, $\operatorname{CSC}(\boldsymbol{V})$ denotes the causal convolution operation on $\boldsymbol{V}$. Let $\boldsymbol{M}_{\operatorname{CSC}}$ represent the sparse matrix associated with the causal convolution and non-linear activation function. $\boldsymbol{M}_{eq}$ denotes the equivalent full attention map. Since $M_{\operatorname{CSC}}$ has the potential to be a full-rank matrix, which increases the upper bound of the rank of the equivalent attention matrix. To this end, we only use a little computational overhead to improve the performance of RELA.
To address the challenge of differentiating between temporary local preferences and genuine shifts in long-term user interests, we design a gated mechanism to intelligently modulate RELA’s output. We utilize a SiLU gate (Qin et al., 2023; Gu and Dao, 2024) and functioning as an output gate, learns to derive weights from the interaction sequence. These weights then adjust the influence of different signals within the user behavior sequence processed by RELA. This allows the model to discern whether a new burst of user behaviors represents a strong but potentially transient local preference or signifies a true evolution in the user’s underlying taste.
The overall structure of the GRELA Block is shown as follows:
(21) $\boldsymbol{H}_{g}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}\right),$
$\boldsymbol{H}_{1}$ $=\operatorname{SiLU}\left(\boldsymbol{W}_{1}\boldsymbol{H}_{g}+\boldsymbol{b}_{1}\right),$
$\boldsymbol{H}_{2}$ $=\operatorname{RELA}\left(\boldsymbol{W}_{2}\boldsymbol{H}_{g}+\boldsymbol{b}_{2}\right),$
where $\boldsymbol{H}_{g}$ is the input of the GRELA block, i.e., when $L=1$, the input $\boldsymbol{H}$ is from the embedding layer, $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{D\times D}$ are the weight matrices of the linear projection, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$ are the biases. Then a Hadamard product is performed on $\boldsymbol{H}_{1}$ and $\boldsymbol{H}_{2}$ followed by a linear projection and Droppath (Larsson et al., 2017) operation:
(22) $\boldsymbol{H}_{3}=\operatorname{DropPath}\left(\boldsymbol{W}_{3}\left(\boldsymbol{H}_{1}\odot\boldsymbol{H}_{2}\right)+\boldsymbol{b}_{3}\right)+\boldsymbol{H},$
where $\boldsymbol{W}_{3}\in\mathbb{R}^{D\times D}$ is the weight matrix, $\boldsymbol{b}_{3}$ is the bias.
Meanwhile, due to the linear projection before the RELA, we can treat it as an input gate of linear attention like long-term memory network (LSTM), previous studies (Greff et al., 2017; Van Der Westhuizen and Lasenby, 2018) have demonstrated that the input gate plays a less significant role compared to the forget gate in LSTM. Inspired by this, we remove the linear projection for the value matrix of linear attention. To this end, we only use the weight matrices on query and key, reducing training costs.
3.3. Multi-Layer Perceptron
To capture the complex features, we further leverage a Multi-Layer Perceptron (MLP) (Kang and McAuley, 2018):
(23) $\boldsymbol{H}_{4}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}_{3}\right),$
$\boldsymbol{H}_{L}$ $=\operatorname{Drop}\left(\operatorname{Drop}\left(\operatorname{GELU}\left(\boldsymbol{H}_{4}\boldsymbol{W}_{4}+\boldsymbol{b}_{4}\right)\right)\boldsymbol{W}_{5}+\boldsymbol{b}_{5}\right)+\boldsymbol{H}_{3},$
where $\boldsymbol{W}_{4}\in\mathbb{R}^{D\times 4D},\boldsymbol{W}_{5}\in\mathbb{R}^{4D\times D}$ denote the weight matrices, $\boldsymbol{b}_{4}$, $\boldsymbol{b}_{5}$ are also the biases. $\boldsymbol{H}_{L}$ indicates the final output representation, and $\boldsymbol{H}_{4}$ denotes the output of layer normalization after the GRELA block. To stabilize the training process and enhance gradient flow through the network, we employ a residual connection (He et al., 2016). For generality, the subscript (0) indicates that the final user representation is obtained using a single GRELA layer.
3.4. Prediction Layer and Training
After the computation of $L$ GRELA blocks and the MLP, we get the final representation of behavior sequences with user interest. Based on this, we compute the dot product between it and the embeddings of all items $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$. This result is then scaled using the Softmax function to obtain the interaction probability score:
(24) $\hat{\mathcal{Y}}=\operatorname{Softmax}(\boldsymbol{E}^{\top}\boldsymbol{H}_{L}).$
Hence, following previous works (Kang and McAuley, 2018; Liu et al., 2024b), we expect to minimize the cross-entropy between the predicted recommendation results $\hat{\mathcal{Y}}$, and the ground truth items $\mathcal{Y}$:
(25) $\mathcal{L}_{CE}\left(\mathcal{Y},\hat{\mathcal{Y}}\right)=\mathcal{Y}\log\left(\hat{\mathcal{Y}}\right)+\left(1-\mathcal{Y}\right)\left(1-\log\left(\hat{\mathcal{Y}}\right)\right).$
3.5. Complexity Analysis
We analyze the computational complexity of RELA (as shown in Eq. (12)), focusing on several key components. To begin with, the kernel function in matrices $\boldsymbol{Q},\boldsymbol{K},\in\mathbb{R}^{N\times{d}}$, incurs a computational cost of $\mathcal{O}(Nd)$. Subsequently, the RoPE operator $\mathcal{R}(\cdot)$ contributes a time complexity that is also $\mathcal{O}(Nd)$. Furthermore, the matrix multiplication operation has a $\mathcal{O}(Nd^{2})$ complexity. This contrasts with the original dot-product attention mechanism, which exhibits a complexity of $\mathcal{O}(N^{2}d)$. Consequently, the overall time complexity of the RELA is $\mathcal{O}(Nd^{2})$.
4. Experiments
Table 3. Statistics of the experimental datasets.
Datasets #Users #Items #Interactions Avg. UA Avg. IA Sparsity
ML-1M 6,041 3,417 999,611 165.5 292.6 95.15%
ML-20M 138,494 18,346 19,984,024 144.3 1,098.3 99.21%
ML-32M 200,949 43,885 31,921,467 158.9 727.4 99.63%
Netflix 120,378 16,087 21,767,054 180.8 1,353.1 98.86%
4.1. Experimental Settings
4.1.1. Datasets
We conduct extensive experiments and verify the effectiveness of our approach on four real-world datasets, including MovieLens-1M, MovieLens-20M, MovieLens-32M  [^2] https://grouplens.org/datasets/movielens/ (dubbed as ML-1M, ML-20M, ML-32M) and Netflix  [^3] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data. The dataset statistics are shown in Table 3, where Avg. UA and Avg. IA denote the average actions of users and items. Notably, for MovieLens datasets, we filter out users and items with fewer than 5 interactions. Netflix is a huge-scale dataset, containing approximately 100 million interactions and five thousand Avg. IA. To simplify the training process, we filter out users with fewer than 50 interactions while retaining items with interactions ranging from 50 to 15,000 for this dataset.
4.1.2. Baseline Models
To verify the effectiveness and efficiency of RecGRELA, we make comparisons between RecGRELA and strong RNN-, Attention-, and Mamba-based baselines:
  • LinRec (Liu et al., 2023): This model reduces the computational costs via linear attention in transformer-based models. We choose SASRec (Kang and McAuley, 2018) as the backbone of LinRec.
  • LRURec (Yue et al., 2024): This method is based on linear recurrent units (Orvieto et al., 2023) for sequential recommendation.
  • Mamba4Rec (Liu et al., 2024b): This is the first Mamba-based method, which explores applying the Mamba architecture in sequential recommendation tasks.
  • EchoMamba4Rec (Wang et al., 2024): This method processes sequences both forward and backward to enhance the sequential representation with Mamba and frequency domain filter.
  • EulerFormer (Tian et al., 2024): This method proposes a transformer variant with complex vectors based on Euler’s formula and a contrastive learning method to improve contextual representations.
  • RecBLR (Liu et al., 2024a): This model shows a behavior-dependent linear recurrent unit to enhance user behavior modeling with a hardware-aware parallel scan approach.
  • SIGMA (Liu et al., 2025): This model employs a flipped Mamba to build a bidirectional framework to improve short sequence representations and contextual modeling.
Table 4. Performance comparisons on ML-1M, ML-20M, ML-32M, and Netflix datasets. Best and second-best results per row are shown in bold and underlined format, respectively. ”Improv.” denotes the percentage improvement of the best-performing model over the second-best for each metric.
Dataset Metric LinRec LRURec Mamba4Rec EchoMamba EulerFormer RecBLR SIGMA RecGRELA Improv.
ML-1M HR@5 0.2187 0.2134 0.2211 0.2243 0.2159 0.2293 0.2325 0.2364 1.68%
NDCG@5 0.1472 0.1418 0.1521 0.1539 0.1493 0.1604 0.1587 0.1653 3.05%
MRR@5 0.1263 0.1217 0.1314 0.1328 0.1281 0.1377 0.1345 0.1419 3.05%
HR@10 0.3109 0.3026 0.3230 0.3253 0.3089 0.3258 0.3291 0.3311 0.61%
NDCG@10 0.1837 0.1734 0.1869 0.1907 0.1768 0.1917 0.1898 0.1959 2.19%
MRR@10 0.1448 0.1340 0.1453 0.1496 0.1365 0.1507 0.1472 0.1545 2.52%
ML-20M HR@5 0.1886 0.1719 0.1813 0.1837 0.1782 0.1890 0.1945 0.2007 3.19%
NDCG@5 0.1297 0.1146 0.1248 0.1271 0.1219 0.1316 0.1361 0.1415 3.97%
MRR@5 0.1103 0.0941 0.1047 0.1072 0.1015 0.1127 0.1169 0.1220 4.36%
HR@10 0.2685 0.2572 0.2649 0.2679 0.2482 0.2664 0.2729 0.2791 2.27%
NDCG@10 0.1554 0.1488 0.1537 0.1578 0.1432 0.1559 0.1613 0.1667 3.35%
MRR@10 0.1209 0.1153 0.1222 0.1247 0.1112 0.1222 0.1272 0.1324 4.09%
ML-32M HR@5 0.1568 0.1513 0.1621 0.1649 0.1587 0.1671 0.1698 0.1802 6.12%
NDCG@5 0.1034 0.0989 0.1092 0.1117 0.1056 0.1151 0.1180 0.1258 6.61%
MRR@5 0.0871 0.0826 0.0923 0.0948 0.0894 0.0980 0.1009 0.1080 7.04%
HR@10 0.2397 0.2235 0.2372 0.2401 0.2208 0.2378 0.2426 0.2519 3.83%
NDCG@10 0.1402 0.1296 0.1369 0.1398 0.1270 0.1379 0.1423 0.1490 4.71%
MRR@10 0.1090 0.1015 0.1062 0.1083 0.1003 0.1073 0.1116 0.1175 5.29%
Netflix HR@5 0.0900 0.0712 0.0809 0.0827 0.0781 0.0873 0.0897 0.0922 2.44%
NDCG@5 0.0606 0.0443 0.0538 0.0559 0.0512 0.0581 0.0604 0.0623 2.81%
MRR@5 0.0507 0.0369 0.0446 0.0467 0.0431 0.0483 0.0508 0.0525 3.35%
HR@10 0.1368 0.1196 0.1305 0.1327 0.1191 0.1338 0.1359 0.1386 1.32%
NDCG@10 0.0765 0.0650 0.0728 0.0736 0.0648 0.0747 0.0763 0.0792 3.53%
MRR@10 0.0582 0.0484 0.0532 0.0553 0.0495 0.0567 0.0582 0.0611 4.98%
Figure 3. RecGRELA vs. baselines in terms of the GPU memory and FLOPs of the training stage on the ML-1M dataset.
4.1.3. Evaluation Protocols
We evaluate performance using three standard metrics: Hit Rate (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). These metrics are assessed based on top-$K$ recommendation lists, specifically using cutoff values $K$ = 5 and $K$ = 10, denoted as HR@5, NDCG@5, MRR@5, HR@10, NDCG@10, and MRR@10, respectively.
4.1.4. Implementation Details
We utilize the RecBole library  [^4] https://github.com/RUCAIBox/RecBole (Zhao et al., 2021; Xu et al., 2023), a unified framework for implementing sequential recommendations. To ensure a fair comparison, we fix the embedding size to 64, the hidden size of the MLP to 256, and the training batch size to 2048 for all the models. We directly adopt the values reported in their original manuscripts for the remaining hyperparameters of the baseline models. The number of blocks $L$ is set to 4. The number of attention heads $h$ is set to 4. The kernel size for 1D causal convolution is 4. We set the Dropout and Droppath rate to 0.1. We use Adam (Kingma, 2014) with a learning rate of 0.001 for optimizing our model. The maximum sequence length $N$ is set to 200 for all the datasets. The early stopping epochs are 20 for ML-1M and 10 for other datasets. All the experiments are conducted on a single NVIDIA RTX4090.
4.2. Overall Performance
Table 4 reports the recommendation performance of RecGRELA compared to seven baselines on four datasets. From the table, we find several observations:
  • Our RecGRELA outperforms all competing transformer-based, RNN-based, and Mamba-based baselines, with improvements ranging from 0.61% to 4.98%. Such a comparison highlights the effectiveness of our design in the long-term sequential recommendation.
  • The performance improvements are more significant on larger datasets with richer interactions and more average action of items. For example, the Netflix dataset has almost 22M interactions and relatively long item actions (1.3K). Conversely, performance gains are more modest for sparser datasets with fewer interactions, users, and items, such as ML-1M. This phenomenon can be attributed to the data-rich environments providing sufficient information for the model to learn more complex patterns, leading to more accurate recommendation performance.
  • The improvements in the metrics of NDCG and MRR are typically greater than those in HR. This discrepancy stems from their distinct evaluation criteria: NDCG and MRR employ position-sensitive weighting schemes prioritizing ranking precision, whereas HR operates as a binary relevance indicator. This ranking superiority enables RecGRELA to generate recommendations with enhanced positional accuracy, elevating user-preferred items to top-tier placements while maintaining contextual relevance. This can also be attributed to introducing RoPE, which enhances the model’s ability to capture positional information.
4.3. Efficiency Analysis
We compare the training efficiency of our RecGRELA with state-of-the-art sequential recommendation models in terms of GPU memory and FLOPs. The results are shown in Fig. 3. The results are based on the ML-1M dataset with a sequence length of 200. Our RecGRELA demonstrates superior efficiency and performance, achieving the highest NDCG@10 of 0.1959 with the lowest GPU memory (12.7GB) among the compared models.
[…]
Table 5. Sensitivity analysis on different numbers of layers. The sequence length is set to 200. Time indicates the training time per epoch, and Memory indicates the GPU memory.
#Layers HR@5 NDCG@5 MRR@5 HR@10 NDCG@10 MRR@10 Time Memory
1 0.1512 0.0987 0.0815 0.2874 0.1669 0.1301 20.70s 5.96GB
2 0.1699 0.1050 0.0899 0.3184 0.1838 0.1426 39.49s 7.71GB
3 0.1753 0.1088 0.0921 0.3228 0.1866 0.1450 58.00s 10.16GB
4 0.1865 0.1152 0.0988 0.3311 0.1959 0.1545 76.50s 12.70GB
5 0.1801 0.1115 0.0954 0.3280 0.1916 0.1499 94.88s 15.14GB
6 0.1776 0.1109 0.0948 0.3242 0.1912 0.1506 113.43s 17.68GB
Figure 4. The effectiveness of each variant of RecGRELA on ML-1M.
4.4. Ablation Study
4.4.1. Core Components
To investigate how different components contribute to our design, we evaluate RecGRELA variants by removing the key components one by one and implementing some modifications as follows:
  • w/ $\boldsymbol{V}$ Matrix, w/o Input Gate: Keep the value $\boldsymbol{V}$ and remove the input linear projection before RELA.
  • w/ $\boldsymbol{V}$ Matrix, w/o SiGLU w/: Remove the SiLU-based gate and keep the value $\boldsymbol{V}$, this variant verifise the effectiveness of the gated mechanism.
  • w/o Droppath: Remove the Droppath operation in the GRELA block.
  • w/o RoPE: Remove the RoPE in the linear attention.
  • w/o Local Shortcut: Remove the Local shortcut from the linear attention.
Fig. 4 records two evaluation metrics, i.e., NDCG@10, and MRR@10 on the ML-1M dataset. From this figure, we have the following observations:
  • The removal of the gating mechanism while retaining the value matrix leads to the most significant performance degradation (e.g., HR@10 drops from 0.3311 to 0.3134 on HR@10), explicitly highlighting the significant role of the SiLU-based gating mechanism in harmonizing long-range sequence dependencies with local preference modeling.
  • Removing the local shortcut also induces substantial performance degradation (e.g., HR@10 drops by 1% compared to the full model), demonstrating that incorporating local preferences in long-sequence recommendation systems significantly enhances overall recommendation quality.
  • The removal of RoPE incurs moderate performance degradation, demonstrating that RoPE enables efficient modeling of relative positional relationships within user interaction sequences.
  • Removing any of the three core contributions in RecGRELA results in substantial performance degradation, particularly in ranking-oriented metrics such as NDCG@10 and MRR@10. This aligns with the observations in Section 4.2.
Figure 5. The performance of different position encodings and activation functions.
4.4.2. Position Encoding and Activation Function
We further investigate the impact of position encoding strategies on RecGRELA. As shown in Fig. 5, the proposed RoPE integration demonstrates superior performance across all metrics while achieving faster training speeds and lower memory consumption than alternative position encoding methods. These empirical results validate the effectiveness of RoPE in enhancing representational capacity within the RecGRELA framework.
We also evaluate the impact of different activation functions on the model’s performance. We replace the function in the gated mechanism and the MLP. As shown in Fig. 5, the SiLU activation function outperforms ReLU and GELU across all metrics, demonstrating the superior performance of SiLU in capturing complex user behavior and enhancing the model’s representational capacity.
Figure 6. Impact of different stacked layers on computational costs. We set the sequence length to 100 to avoid GPU memory errors.
4.5. Sensitivity Analysis
In this section, we conduct experiments on the ML-1M dataset to analyze the influence of three hyperparameters: the sequence length $N$, the number of stacked GRELA blocks $L$, the number of attention heads $h$, and the Dropout rates.
4.5.1. The Number of Layers
The results in Table 5 reveal that stacking more RecGRELA layers fails to consistently enhance recommendation quality while markedly degrading computational efficiency, primarily due to the overfitting problem of multiple RecGRELA layers. Specifically, model performance peaks at four stacked layers, beyond which further depth expansion leads to metric degradation with substantial computational overhead. Notably, the three-layer configuration achieves competitive performance with state-of-the-art baselines, demonstrating the architecture’s effectiveness in balancing representational capacity and operational efficiency. Furthermore, our comparative analysis of layer-wise efficiency between RecGRELA and SIGMA reveals distinct scaling patterns. As illustrated in Fig. 6, the growth trend of computational occupation of RecGRELA is relatively gradual, whereas that of SIGMA shows a steeper increase, indicating that RecGRELA exhibits better scaling capability when the training requires more layers.
Figure 7. Impact of sequence length and attention head variations on computational costs. When investigating the impact of sequence length, we set the attention head to 2 for EulerFormer and 4 for RecGRELA. When investigating the impact of attention heads, we set the sequence length to 100 to avoid GPU memory errors. Other hyperparameters are the same as the main experiments.
4.5.2. The Number of Sequence Length
To examine training cost variations across sequence lengths in long-term SRSs, we evaluate maximum sequence lengths from 100 to 200 with 25-step increments. Fig. 7 demonstrates that RecGRELA maintains consistently lower computational complexity than EulerFormer across all tested configurations.
To evaluate the model’s performance characteristics with varying input lengths, we varied the maximum sequence length processed by the model, testing values from 100 to 200, while keeping other hyperparameters constant. The impact on recommendation performance was measured using NDCG@10 and MRR@10, as illustrated in Fig. 8. The green and blue lines show the model’s performance exhibits relative stability for sequence lengths between 100 and 150. Notably, performance significantly improves when the sequence length increases to 200. This suggests that while the model performs consistently well with shorter sequences, capturing longer-range dependencies present in sequences of length 200 yields a clear benefit in terms of recommendation performance.
4.5.3. The Number of Attention Heads
We evaluate the impact of the number of attention heads on model efficiency. As shown in Fig. 7, the results demonstrate that RecGRELA maintains stable GPU memory and training time across varying head counts, while EulerFormer exhibits proportional increases in computational overhead with expanded attention heads. RecGRELA applies rotary position encoding directly to the full embedding dimension, avoiding repetitive computations across attention heads as EulerFormer does. Our method proves that handling position encodings at the embedding level saves computation compared to per-head processing.
We further investigated the model’s performance with respect to the number of attention heads. The results are also shown in Fig. 8. Both metrics consistently demonstrate that using four attention heads yields the best performance for our model configuration on this dataset. Utilizing fewer heads (two) appears insufficient to capture the complexity of user-item interactions while employing significantly more heads does not offer further benefits and may even slightly degrade performance. This could be attributed to potential redundancy or fragmentation of attention patterns with excessive heads. Therefore, four attention heads represent the optimal setting identified in our experiments.
4.5.4. The Dropout and Droppath Rates
Regularization techniques like Dropout and Droppath prevent overfitting and improve model generalization. We assessed the model’s sensitivity to the rates of these two techniques, varying each from 0.1 to 0.5. Fig. 8 displays the impact on NDCG@10 and MRR@10. The sensitivity analysis suggests that lower regularization rates (0.1 for both) are preferable. Higher rates, particularly for Dropout, lead to substantial performance degradation. This highlights the importance of carefully tuning regularization hyperparameters, as overly aggressive regularization can harm performance more than it helps prevent overfitting in sequential recommendation.
Figure 8. Influence of the sequence length, head numbers, Dropout, and Droppath on various metrics.
4.6. Visualization Analysis
To further investigate the impact of the proposed Local Shortcut on the representational capacity of our model, particularly regarding the theoretical motivation to increase feature matrix rank, we conduct a visualization analysis on the ML-1M dataset. We visualize the output feature maps from the last layer for three distinct models: first, LinRec, which serves as a standard linear attention baseline; second, GRELA without the Local Shortcut, representing our architecture with this specific component ablated; and third, the complete GRELA model, labeled as Ours. The resulting heatmaps display the feature matrix structure for a sequence length $N$ of 200 and a feature dimension $d$ of 64. We also compute the rank for each matrix. Given the feature dimension $d$ = 64, this is the maximum possible rank in this configuration.
Fig. 9 helps show why our design is effective by visualizing the resulting feature matrices. The heatmap on the left (LinRec), shows a feature matrix with a rank of 47. The middle heatmap is for GRELA without the Local Shortcut, yielding a feature matrix with the lowest rank at just 32. Visually, strong horizontal lines and repeating patterns indicate that the features are too similar and lack local detail. This low matrix rank confirms reduced expressive power and information content. In contrast, the heatmap on the right, from our full GRELA model with the Local Shortcut, displays a much more complex and detailed feature matrix. The strong lines disappear, showing more varied features. Significantly, its matrix rank increases substantially to 58, closely approaching the theoretical maximum rank of 64.
The Local Shortcut, which uses causal convolution, adds important local details to the features. This helps fix the low-matrix rank problem that happens with just linear attention. Adding these local details makes the overall matrix rank higher. As a result, the features become richer and show more variety. We can see this improvement in the heatmap, and the higher matrix rank confirms it. This proves the Local Shortcut works well to get around the limits of linear attention. It helps the model capture immediate local preferences and crucial long-term dependencies within the user’s interaction sequence.
Figure 9. Heatmap of output feature maps from the last layer.
5. Related Work
5.1. Sequential Recommendation
Earlier fundamental models of sequential recommendation focus on using Convolutional Neural Networks (Tang and Wang, 2018; Yuan et al., 2020) and Recurrent Neural Networks (RNNs) (Hidasi, 2016; Li et al., 2017). Due to RNNs’ performance limits and poor parallel processing, recent sequential recommendation studies have increasingly turned to Transformer-based architectures (Wu et al., 2020; Li et al., 2021; Du et al., 2023; Fan et al., 2022; Zhang et al., 2023; Zhao et al., 2023; Li et al., 2023a; Shi et al., 2024) or MLPs (Long et al., 2024; Zhou et al., 2022; Li et al., 2023b, 2022; Gao et al., 2024; Jiang et al., 2024) for better efficiency and scalability. The most representative work is SASRec (Kang and McAuley, 2018), which is the first to use the self-attention mechanism for capturing user interest features from interaction sequences. BERT4Rec (Sun et al., 2019) adopts a deep encoder-based bidirectional model with self-attention to model user behavior sequences, achieving strong recommendation performance. However, the Transformer architecture suffers from quadratic computational complexity with respect to sequence length due to the dot-product attention operation, which limits its efficiency for sequence modeling tasks. Recently, many models with linear computational complexity have emerged as alternatives to the original Transformer architecture, such as linear attention (Liu et al., 2023; Tian et al., 2024), linear RNNs (Liu et al., 2024a; Yue et al., 2024), and state space models (Liu et al., 2024b; Qu et al., 2024; Fan et al., 2024). While LinRec (Liu et al., 2023) adopts linear attention to reduce the computational complexity of the dot-product attention, they still rely on learnable position encodings to capture positional information. EulerFormer (Tian et al., 2024) introduces a novel Euler transformation with a rotary form for unified modeling of both semantic and positional difference, yet uses the dot-product attention, which results in computational costs comparable to LinRec. To overcome these specific limitations, RecGRELA combines linear attention with rotary position encodings and introduces a gated mechanism to model both local and long-term user interests.
5.2. Linear Attention
As opposed to dot-product attention, linear attention (Katharopoulos et al., 2020) replaces the Softmax with separate kernel functions, thereby reducing computational complexity to $\mathcal{O}(N)$ through a change in computation order. Though efficient, designing a linear attention module as effective as softmax attention is a nontrivial problem. Although linear attention has achieved some success in other domains, such as computer vision (Shen et al., 2021; Qin et al., 2022; Han et al., 2024a, b) and neural language processing (Wang et al., 2020; Qin et al., 2023; Yang et al., 2024). In sequential recommendation, despite existing linear attention-based methods (Fan et al., 2021; Liu et al., 2023) decreasing the computational complexity, they cannot effectively model position encoding within behavior sequence and fail to capture local preference in long-term interactions. This motivates us to design a more efficient and effective linear attention-based module to address these issues. In response, we present RecGRELA, which aims to realize the full potential of linear attention in sequential recommendation by overcoming these critical challenges in positional awareness and local preference modeling.
6. Conclusion and Future Work
In this work, we propose RecGRELA, a new paradigm for long-term sequential recommendation. By integrating rotary position encoding into a linear attention mechanism, our Rotary-Enhanced Linear Attention module captures long-range dependencies with reduced computational complexity. To model local interaction in long-range dependencies, we design a Local Shortcut operation and provide theoretical insight from the perspective of the rank of the matrix. To address the difficulty of the model in perceiving and adapting to dynamic user preference shifts, we propose a SiLU-based gating mechanism that dynamically balances local and long-term representations. Extensive experiments on four benchmark datasets show that RecGRELA outperforms state-of-the-art methods based on RNNs, Transformers, and Mamba architectures, achieving superior performance with significantly lower memory overhead.
In the future, we plan to explore the application of RecGRELA in other recommendation scenarios, such as session-based and multi-modal recommendations. Since linear attention shows great potential in various domains, we also plan to explore more efficient attention mechanisms for sequential recommendation tasks.
Acknowledgements.
This work is funded by the National Natural Science Foundation of China (Grant No. 72074036, 62072060) and supported by the Natural Science Foundation of Chongqing, China, No. CSTB2024NSCQ-MSX0701, CSTB2024NSCQ-MSX0617.

### v2

@@
-Gated Rotary-Enhanced Linear Attention for Long-term Sequential Recommendation
-Abstract.
-In Sequential Recommendation Systems (SRSs), Transformer models show remarkable performance but face computation cost challenges when modeling long-term user behavior sequences due to the quadratic complexity of the dot-product attention mechanism.
-By approximating the dot-product attention, linear attention provides an efficient option with linear complexity.
-However, existing linear attention methods face two limitations: 1) they often use learnable position encodings, which incur extra computational costs in long-term sequence scenarios, and 2) they may not consider the user’s fine-grained local preferences and confuse these with the actual change of long-term interests.
+Methodology, Writing – original draft, Investigation, Software
+Abstract
+In Sequential Recommendation Systems (SRSs), Transformer models have demonstrated remarkable performance but face computational and memory cost challenges, especially when modeling long-term user behavior sequences.
+Due to its quadratic complexity, the dot-product attention mechanism in Transformers becomes expensive for processing long sequences.
+By approximating the dot-product attention using elaborate mapping functions, linear attention provides a more efficient option with linear complexity.
+However, existing linear attention methods face three limitations: 1) they often use learnable position encodings, which incur extra computational costs in long-term sequence scenarios, 2) they may not sufficiently account for user’s fine-grained local preferences (short-lived burst of interest), and 3) they try to capture some temporary activities, but often confuse these with stable and long-term interests.
+This can result in unclear or less effective recommendations.
 To remedy these drawbacks, we propose a long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention (RecGRELA).
-Specifically, we first propose a Rotary-Enhanced Linear Attention (RELA) module to model long-range dependency within the user’s historical information using rotary position encodings.
-We then introduce a local short operation to incorporate local preferences and demonstrate the theoretical insight.
-We further introduce a SiLU-based Gated mechanism for RELA (GRELA) to help the model determine whether a user’s behavior indicates local interest or a genuine shift in long-term preferences.
-Experimental results on four public datasets demonstrate that our RecGRELA achieves state-of-the-art performance compared to existing SRSs while maintaining low memory overhead.
-Sequential Recommendation, Linear Attention, Rotary Position Encoding, Gated Mechanism
-ccs: Information systems Recommender systems
-1. Introduction
-Figure 1.
-Illustration of differences of SASRec, LinRec, and our RecGRELA.
-$\phi(\cdot)$ denotes the kernel function.
-(a) SASRec utilizes Softmax attention, where similarities between all query-key pairs are computed, leading to quadratic complexity for sequence length.
-(b) LinRec implements linear attention by transforming Q and K with $\phi(\cdot)$.
-This reordering of computations achieves linear complexity, though it may diminish the model’s expressive power.
-(c) Our RecGRELA enhances linear attention by incorporating Rotary Positional Encoding (RoPE) for Q and K and adding a shortcut connection, aiming to achieve strong expressive capability while maintaining linear computational complexity and removing the need for explicit Learned Position Encodings (LPE).
-Sequential Recommender Systems (SRSs) (Wang et al., 2019; Fang et al., 2020; Pan et al., 2024; Niu et al., 2025; Zhang et al., 2025; Yang et al., 2023; Zhang et al., 2021) have become a fundamental technology in personalized services, significantly transforming how online platforms predict user behavior.
+Specifically, we first propose a Rotary-Enhanced Linear Attention (RELA) module to efficiently model long-range dependency within the user’s historical information using rotary position encodings.
+Then, we introduce a local short operation to add the local preferences of interactions and show the theoretical insight.
+We further introduce a SiLU-based Gated mechanism for RELA (GRELA) to help the model tell if a user behavior shows a short-term, local interest or a real change in their long-term tastes.
+Experimental results on four public benchmark datasets show that our RecGRELA achieves state-of-the-art performance compared with existing SRSs based on Recurrent Neural Networks, Transformer, and Mamba while keeping low memory overhead.
+Our source code is available at https://anonymous.4open.science/r/RecGLARE-B4AC.
+keywords:
+[…]
+{highlights}
+Rotary Enhanced Linear Attention integrates RoPE for efficient long-sequence modeling.
+Local Shortcut operation injects dense local user preferences via causal convolution.
+RecGRELA achieves competitive performance with low memory overhead.
+1 Introduction
+Figure 1: Illustration of differences of SASRec, LinRec, and our RecGRELA. $\phi$ denotes the kernel function.
+Sequential Recommender Systems (SRSs) [54, 9, 37, 35, 68, 48] have become a fundamental technology in personalized services, significantly transforming how online platforms predict user behavior.
 By analyzing sequences of user interactions across e-commerce, streaming services, and social media, SRSs capture both short-term interests and long-term preference trends.
@@
 To address the challenge of modeling user behavior sequences, researchers explore various methodologies.
-Traditional approaches, such as those based on Markov Chains (He et al., 2017; Rendle et al., 2010) or Recurrent Neural Networks (RNNs) (Chen et al., 2018; Cui et al., 2018; Donkers et al., 2017; Xu et al., 2019), lay the groundwork but also face limitations in capturing long-range dependencies or suffered from computational inefficiencies during training and inference, particularly with increasing sequence lengths.
-Consequently, the field moves towards more advanced models that can handle complex sequential dependencies.
-This also presents several new challenges for the field.
+Traditional approaches, such as those based on Markov Chains [17, 42] or Recurrent Neural Networks (RNNs) [1, 3, 4, 58], lay the groundwork but also face limitations in capturing long-range dependencies or suffered from computational inefficiencies during training and inference, particularly with increasing sequence lengths.
+Consequently, the field moves towards more advanced models that can handle complex sequential dependencies more effectively.
+This also presents new challenges.
 Specifically, the first challenge is Efficiency.
-Transformer models (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019) emerge as a compelling replacement for earlier approaches like RNNs, leveraging their ability to process sequential data and capture both local and global preferences of users.
-However, the dot-product attention (as show in Fig.
-1(a)) in Transformers exhibits quadratic complexity relative to sequence length, resulting in high costs for long sequences.
-To address this issue, linear attention (Katharopoulos et al., 2020; Liu et al., 2023) has been proposed as an efficient alternative, providing linear complexity (as show in Fig.
-1(b)).
-However, compared to dot-product attention, current linear attention approaches may suffer from performance drops.
-In addition to the efficiency of attention mechanisms, effectively capturing the positional information in user interaction sequences is another crucial challenge that affects both model performance and operational efficiency.
+Transformer models [19, 47, 63] emerge as a compelling replacement for earlier approaches like RNNs, leveraging their ability to process sequential data and capture both local and global preferences of users.
+However, the dot-product attention mechanism in Transformers exhibits quadratic complexity relative to sequence length, resulting in high costs for long sequences.
+To address this issue, linear attention [20, 31] has been proposed as an efficient alternative, providing linear complexity while maintaining the modeling capabilities.
 Position encoding is widely adopted to capture the sequential order of items users interact with due to its ability to enhance the positional awareness of contextual vectors.
 As illustrated in Fig.
-1(a) and (b), the SASRec (Kang and McAuley, 2018) and LinRec (Liu et al., 2023) both rely on learnable position encoding.
+1(a) and (b), the above attention-based methods rely on learnable position encoding.
 The learned positional encoding requires more parameters than fixed positional encoding schemes since each position needs its own learned vector, and they do not have any inherent structure that captures the notion of relative distances between positions.
-To handle these drawbacks, some works (Su et al., 2024; Tian et al., 2024) employ an improved Rotary Position Encoding (RoPE), which injects relative positional information in a more efficient manner.
+To handle these drawbacks, some works [46, 50] employ an improved Rotary Position Encoding (RoPE), which injects relative positional information in a more efficient manner.
 However, they still use the dot-product attention mechanism, which limits its scalability for long sequences.
 Furthermore, the RoPE exhibits inefficiency in multi-head attention scenarios, restricting the model’s overall expressive capability.
-Linear attention, while effective for grasping overarching long-term dependencies, local preferences are equally vital for the long-term sequential recommendation, driven by users’ immediate preferences.
+Linear attention, while effective for grasping overarching long-term dependencies, can struggle to identify Fine-Grained Local Preferences.
+In recommendation contexts, user interactions often reveal dense, localized activity sequences.
 For instance, a user might intensely engage with a specific series of related items.
@@
 Subsequently, they might delve into documentaries that explore the creators behind such cinematic universes.
-However, linear attention’s tendency to smooth features can obscure important local information, while RoPE’s distance-based decay (Su et al., 2024) weakens attention to nearby items, particularly in long sequences.
-Both these characteristics hinder the precise modeling of local patterns.
+These closely related, short-term sequences of engagement are crucial for understanding immediate user intent, yet many existing approaches overlook their significance in long-term sequential recommendation.
 Even when linear attention successfully captures both long-term interests and short-term local preferences, a significant challenge persists: when a user’s preferences really change, it’s hard for the model to identify them.
@@
 Specifically, we design a Rotary Enhanced Linear Attention (RELA) module.
-RELA builds upon linear attention, efficiently solving the computation limitation with a linear complexity of $\mathcal{O}(N)$, It also enhances its positional awareness by incorporating Rotary Position Encoding (RoPE).
-RoPE enables the model to capture relative positional relationships between items in a parameter-efficient manner in multi-head scenarios, overcoming the drawback of learnable positional encodings that introduce additional parameters.
-Concurrently, we design a Local Shortcut operation with causal 1D convolution in RELA to integrate local preferences and present theoretical insights based on the matrix rank analysis.
+RELA builds upon the efficiency of linear attention, directly addressing the quadratic complexity issue of standard Transformers, but significantly enhances its positional awareness by incorporating Rotary Position Encoding (RoPE).
+This allows the model to effectively capture relative inter-item positional relationships in a parameter-efficient manner, overcoming the drawbacks of purely learnable or fixed encodings and the inefficiency of RoPE in standard multi-head dot-product attention.
+Concurrently, it integrates local preferences efficiently via a Local Shortcut operation with causal 1D convolution, and we present theoretical insights based on matrix rank analysis.
 We further introduce a SiLU-based gated mechanism to identify whether user behavior is a temporary burst of interest or a true shift of long-term interest via adjusting the output of RELA.
@@
 By replacing dot-product attention with linear attention, our approach reduces computational costs while maintaining the ability to capture inter-item positional relationships.
-  • We further develop a Local Shortcut operation to inject dense local interaction into the linear attention framework.
+  • We develop a Local Shortcut operation to inject dense local interaction into the linear attention framework.
 This mechanism combines causal 1D convolution, explicitly modeling local preferences of user behaviors without sacrificing computational efficiency.
   • We introduce a SiLU-based gated linear attention mechanism to distinguish true shifts in user preferences from transient local interests.
-  • Based on these designs, we further construct the RecGRELA, a novel long-term sequential recommendation paradigm that achieves competitive performance compared with state-of-the-art SRSs on four real-world datasets, keeping low memory overhead.
-Our source code is available at Github111 https://github.com/TBI805/RecGRELA..
-Table 1. Notations and their descriptions for Section 2.
-Notations Description
-$\mathcal{U}$ user set with each user $u\in\mathcal{I}$
-$\mathcal{V}$ user set with each user $v\in\mathcal{V}$
-$\mathcal{S}_{u}$ interaction sequence for user $u$
-$N$ sequence length
-$D,d$ embedding dimension and head dimension of training sequence
-$\boldsymbol{X}$ input sequence
-$\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}$ projection matrix for $\boldsymbol{X}$
-$\boldsymbol{Q}_{m},\boldsymbol{K}_{m},\boldsymbol{V}_{m}$ query, key and value matrix on position $m$
-$\boldsymbol{O}_{m}$ attention result on position $m$
-$\mathcal{O}\left(\cdot\right)$ computation complexity
-$\phi\left(\cdot\right)$ kernel function
-$\mathbf{R}$ the rotation matrix
-$\mathcal{C}\left(\cdot\right)$, $\mathcal{C}^{-1}\left(\cdot\right)$ operator for converting between real and complex vector representations
-$\mathcal{R}\left(\cdot\right)$ RoPE operation
-$\mathrm{Re}[\cdot]$ real part of the complex vector
-2. PRELIMINARY
-2.1. Problem Definition
-[…]
-2.2. Dot-Product Attention and Linear Attention
-Let $\boldsymbol{X}\in\mathbb{R}^{N\times D}$ denote an input sequence.
+Based on these designs, we further construct the RecGRELA, a novel long-term sequential recommendation paradigm that achieves competitive performance compared with state-of-the-art SRSs on four benchmark datasets, keeping low memory overhead.
+2 PRELIMINARY
+2.1 Problem Definition
+Considering a set of users $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$ and a set of items $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$.
+We use $\mathcal{S}_{u}=\left\{v_{t}\right\}_{t=1}^{T}$ to denote an interaction sequence of $T$ chronologically ordered items that the user $u$ has interacted with.
+Given a user interaction sequence $\mathcal{S}_{u}$, the task of sequential recommendation is to compute a ranking list consisting of Top-$k$ items that the user $u$ is likely to interact with at the next time step $T+1$.
+We define $N$ as the sequential length and $D$ as its corresponding dimension.
+Typically, the $N$ is often relatively short, but in some complex scenarios, where the $N$ satisfies $N$ > 1.5DD (e.g., when DD=64, the $N$ > 96), the sequence is regarded as “long-term[31].
+2.2 Dot-Product Attention and Linear Attention
+Let $\boldsymbol{X}\in\mathbb{R}^{N\times D}$ denotes an input sequence.
 Formally, the input sequence $\boldsymbol{X}$ is projected by three projection matrices $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}\in\mathbb{R}^{D\times d}$, to corresponding representations $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}_{Q}$, $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}_{K}$, and $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}_{V}$.
-The dot-product attention (Vaswani et al., 2017) can be written as:
-(1) $\quad{\boldsymbol{O}}_{m}=\sum_{n=1}^{N}\frac{\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}{\sum_{n=1}^{N}\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}\boldsymbol{V}_{n},$
-where $\boldsymbol{Q}_{m},\boldsymbol{K}_{n},\boldsymbol{V}_{n}\in\mathbb{R}^{1\times d}$, and $d$ is the head dimension.
+The dot-product attention [52] can be written as:
+$\quad{\boldsymbol{O}}_{m}=\sum_{n=1}^{N}\frac{\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}{\sum_{n=1}^{N}\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}\boldsymbol{V}_{n},$ (1)
+where $\boldsymbol{Q}_{m},\boldsymbol{K}_{n},\boldsymbol{V}_{n}\in\mathbb{R}^{1\times d}$, and $d$ is head dimension.
 In this case, calculating attention weights in dot-product attention leads to the computation complexity of $\mathcal{O}(N^{2})$ with respect to the input sequence length $N$.
 For autoregressive modeling, dot-product attention can be implemented by incorporating a masking mechanism, which restricts the influence on the $m$-th position to a position $n$ satisfying $n\leq m$.
-Linear attention (Katharopoulos et al., 2020) is proposed as an effective alternative that reduces the computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ compared with dot-product attention.
+Linear attention [20] is proposed as an effective alternative that reduces the computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ compared with dot-product attention.
 Specifically, linear attention modifies the dot-product attention by replacing the non-linear Softmax operation with linear normalization, while incorporating an additional kernel function $\phi$ in query matrix $\boldsymbol{Q}$ and key matrix $\boldsymbol{K}$:
-(2) $\boldsymbol{O}_{m}=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}{\sum_{n=1}^{N}\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}\boldsymbol{V}_{n}.$
+$\boldsymbol{O}_{m}=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}{\sum_{n=1}^{N}\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}\boldsymbol{V}_{n}.$ (2)
 By utilizing the associative property of matrix multiplication, the computation is reordered from $(\boldsymbol{Q}\boldsymbol{K}^{\top})\boldsymbol{V}$ to $\boldsymbol{Q}(\boldsymbol{K}^{\top}\boldsymbol{V})$:
-(3) $\boldsymbol{O}_{m}=\frac{\phi\left(\boldsymbol{Q}_{m}\right)\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
+$\boldsymbol{O}_{m}=\frac{\phi\left(\boldsymbol{Q}_{m}\right)\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$ (3)
 where the computational complexity is decreased to $\mathcal{O}(N)$.
-In (Katharopoulos et al., 2020), the kernel function employs the Exponential Linear Unit (ELU) activation (Clevert, 2015), which prevents gradient vanishing when processing negative values of $\boldsymbol{X}$.
+In [20], the kernel function employs the Exponential Linear Unit (ELU) activation [2], which prevents gradient vanishing when processing negative values of $\boldsymbol{X}$.
 In our work, we follow this setting for a fair comparison.
-2.3. Position Encoding
-Absolute Position Encoding (APE) (Vaswani et al., 2017) incorporates fixed values into the embeddings of sequence elements, establishing a consistent reference frame across all tokens.
+2.3 Position Encoding
+Absolute Position Encoding (APE) [52] incorporates fixed values into the embeddings of sequence elements, establishing a consistent reference frame across all tokens.
 For a token at the position $t$ within a sequence of length $N$, this encoding is computed through a deterministic function utilizing paired sine and cosine operations across different frequencies.
 The formulation is denoted as follows:
-(4) $PE\left(p,2t\right)$ $=\sin\left(p/10000^{2t/\left(D/2\right)}\right),$
+$PE\left(p,2t\right)$ $=\sin\left(p/10000^{2t/\left(D/2\right)}\right),$ (4)
 $PE\left(p,2t+1\right)$ $=\cos\left(p/10000^{2t/\left(D/2\right)}\right),$
 where p is the position of the item, t is the dimension of p.
-Learnable Position Encoding (LPE) (Wang et al., 2021; Kang and McAuley, 2018) adopts a distinct methodology by modeling positional information as learnable parameters that are optimized during the training process.
+Learnable Position Encoding (LPE) [53, 19] adopts a distinct methodology by modeling positional information as learnable parameters that are optimized during the training process.
 Specifically, for an item embedding x at position $m$, the transformation is performed as follows:
-(5) $\textbf{{x}}_{m}=\textbf{{x}}+p_{m},$
+$\textbf{{x}}_{m}=\textbf{{x}}+p_{m},$ (5)
 where $p_{i}$ denotes the APE or LPE.
-Rotary Position Encoding (RoPE) (Su et al., 2024) captures relative positional information (Shaw et al., 2018) through absolute position encoding methods.
-To achieve this, RoPE encodes the relative positional information via a rotation matrix $\mathbf{R}\in\mathbb{C}^{N\times(d/2)}$:
-(6) $\mathbf{R}\left(m,t\right)=e^{i\theta_{t}m},$
-[…]
-(7) $\bar{\boldsymbol{Q}}^{\prime}_{m}$ $=\boldsymbol{\bar{Q}}_{m}\odot\mathbf{R}\left(m,t\right),$
+Rotary Position Encoding (RoPE) [46] captures relative positional information [43] through absolute position encoding methods via encoding the relative positional information with a rotation matrix $\mathbf{R}\in\mathbb{C}^{N\times(d/2)}$:
+$\mathbf{R}\left(m,t\right)=e^{i\theta_{t}m},$ (6)
+where $i$ denotes the imaginary unit in complex numbers.
+Following [52], the $\theta_{t}$ is set to $10000^{-t/(d/2)}$, where $t\in[1,2,...,d/2]$.
+Then RoPE converts $\boldsymbol{Q}_{m}$, $\boldsymbol{K}_{n}\in\mathbb{R}^{N\times{d}}$ to complex vector $\boldsymbol{\bar{Q}}_{m}$, $\boldsymbol{\bar{K}}_{n}\in\mathbb{C}^{N\times(d/2)}$ by considering (2tt)-th dim as real part and (2tt + 1)-th dim as imaginary part.
+The RoPE is applied to complex vectors of query and key as:
+$\bar{\boldsymbol{Q}}^{\prime}_{m}$ $=\boldsymbol{\bar{Q}}_{m}\odot\mathbf{R}\left(m,t\right),$ (7)
 $\bar{\boldsymbol{K}}^{\prime}_{n}$ $=\boldsymbol{\bar{K}}_{n}\odot\mathbf{R}\left(n,t\right),$
@@
 We can define the RoPE operation $\mathcal{R}\left(\cdot\right)$ to an input real vector $\boldsymbol{X}$ as follows:
-(8) $\mathcal{R}\left(\boldsymbol{X}\right)=\mathcal{C}^{-1}\left(\mathcal{C}\left(\boldsymbol{X}\right)\odot\mathbf{R}\right),$
+$\mathcal{R}\left(\boldsymbol{X}\right)=\mathcal{C}^{-1}\left(\mathcal{C}\left(\boldsymbol{X}\right)\odot\mathbf{R}\right),$ (8)
 where $\mathcal{C}\left(\cdot\right)$ converts a real-valued tensor into a complex-valued tensor, and $\mathcal{C}^{-1}\left(\cdot\right)$ converts a complex-valued tensor into a real-valued tensor.
-Table 2. Notations and their descriptions for Section 3.
-Notations Description
-$\boldsymbol{E}$ learnable embedding matrix
-$\boldsymbol{H}_{o}$, $\boldsymbol{H}$ initial and final embedding matrix
-$\boldsymbol{H}_{g}$ input of the the GRELA block
-$\boldsymbol{H}_{L}$ final output for prediction
-$\boldsymbol{O}_{f}$ output of the RELA
-$\boldsymbol{S}$ output of the causal convolution
-$\operatorname{CSC}(\cdot)$ causal convolution operation
-$\boldsymbol{M}_{\operatorname{CSC}}$ sparse matrix associated with causal convolution and non-linear activation function
-$\boldsymbol{M}_{eq}$ equivalent full attention map
-$\boldsymbol{W}_{1}$, …, $\boldsymbol{W}_{5}$ weight matrix for linear projection
-$\boldsymbol{b}_{1},...,\boldsymbol{b}_{5}$ bias for linear projection
-$\boldsymbol{O}_{i}$ attention result on position $i$
-$\hat{\mathcal{Y}}$ predicted recommendation result
-$\mathcal{L}_{CE}$ cross-entropy loss
-3. METHODOLOGY
-In this section, we introduce our proposed framework, RecGRELA. The framework overview is shown in Figure 2.
-Figure 2.
-The overview of our proposed architecture.
-The Overall Architecture (left) processes an input sequence through an Embedding Layer, which is then fed into a stack of $L$ identical encoder layers.
+3 METHODOLOGY
+Figure 2: The overview of our proposed architecture.
+The Overall Architecture (a) processes an input sequence through an Embedding Layer, which is then fed into a stack of $L$ identical encoder layers.
 Each encoder layer consists of Layer Normalization, a GRELA Block, another Layer Normalization, and an MLP, finally leading to a Prediction Layer.
-The central components are the GRELA Block (middle-bottom), which utilizes RoPE (Rotary Position Encoding, middle-top), and the RELA module (right).
+The central components are the GRELA Block (b), which utilizes RoPE (Rotary Position Encoding, d), and the RELA module (c).
 Within the RELA module, query ($Q$) and key ($K$) are processed by ELU activation and the RoPE.
@@
 A Causal Conv1D and SiLU activation processes the value ($V$) projection to capture local context.
+In this section, we introduce our proposed framework, RecGRELA.
+The framework overview is shown in Figure 2.
 The presentation is structured as follows: initially, we introduce the embedding layer; secondly, we elaborate on the core components of the GRELA Block; subsequently, we present the details of the Multi-Layer Perceptron, and finally, we describe the prediction layer and training strategies.
-The key notations of this section are listed in Table 2.
-3.1. Embedding Layer
-Following (Kang and McAuley, 2018), we utilize an embedding layer to map the training sequence $\mathcal{S}_{u}$ into a high-dimensional space.
+3.1 Embedding Layer
+Following [19], we utilize an embedding layer to map the training sequence $\mathcal{S}_{u}$ into a high-dimensional space.
 The embedding layer uses a learnable embedding matrix $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $\mathcal{V}$ denotes the item set and $D$ represents the embedding dimension.
 Given an input interaction sequence $\mathcal{S}_{u}$ of length $N$, it can be projected into the initial embedding matrix $\boldsymbol{H}_{o}\in\mathbb{R}^{N\times D}$ as follows:
-(9) $\boldsymbol{H}_{o}=\boldsymbol{E}\mathcal{S}_{u}.$
-Then we apply the Dropout (Srivastava et al., 2014) and layer normalization (Lei Ba et al., 2016) operations to obtain the final embedding $\boldsymbol{H}\in\mathbb{R}^{N\times D}$:
-(10) $\boldsymbol{H}=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\boldsymbol{H}_{o}\right)\right).$
-Attention-based sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020; Liu et al., 2023) typically inject a learnable position encoding into the final embedding, while our method captures position awareness with rotary position encoding by multiplying the rotation matrix with outputs of the kernel function in linear attention.
+$\boldsymbol{H}_{o}=\boldsymbol{E}\mathcal{S}_{u}.$ (9)
+Then we apply the Dropout [45] and layer normalization [23] operations to obtain the final embedding $\boldsymbol{H}\in\mathbb{R}^{N\times D}$:
+$\boldsymbol{H}=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\boldsymbol{H}_{o}\right)\right).$ (10)
+Attention-based sequential recommendation models [19, 47, 26, 31] typically inject a learnable position encoding into the final embedding, while our method captures position awareness with rotary position encoding by multiplying the rotation matrix with outputs of the kernel function in linear attention.
 This design eliminates the need to incorporate positional information into the embedding layers.
-3.2. GRELA Block
+3.2 GRELA Block
 In this part, we present the design of our proposed GRELA Block, which includes the Rotary-Enhanced Linear Attention module and the Local Shortcut operation.
-3.2.1. Rotary-Enhanced Linear Attention
+3.2.1 Rotary-Enhanced Linear Attention
 Overall Design.
-Previous work (Liu et al., 2023) provides $\mathcal{O}(N)$ computational complexity of attention mechanism to alleviate the issue of quadratic complexity in dot-product attention for sequential recommendation.
+Previous work [31] provides $\mathcal{O}(N)$ computational complexity of attention mechanism to alleviate the issue of quadratic complexity in dot-product attention for sequential recommendation.
 However, it ignores the high cost of learnable position encoding in long-sequence modeling.
-Meanwhile, some works (Tian et al., 2024; Lopez-Avila et al., 2024) notice these issues and propose improved RoPE to address them, but unfortunately, they fail to address the problem of dot-product attention with $\mathcal{O}(N^{2})$ complexity.
+Meanwhile, some works [50, 34] notice these issues and propose improved RoPE to address them, but unfortunately, they fail to address the problem of dot-product attention with $\mathcal{O}(N^{2})$ complexity.
 To bridge this gap, we propose a novel Rotary-Enhanced Linear Attention (RELA) mechanism, aiming to simultaneously resolve the issues of quadratic complexity and ineffective position encoding in long-sequence modeling.
@@
 (3) as follows:
-(11) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
+$\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$ (11)
 where $\mathcal{R}\left(\cdot\right)$ denotes injecting the RoPE into queries or keys.
-Following (Katharopoulos et al., 2020), we set $\phi(x)=\operatorname{ELU}(x)+1$.
+Following [20], we set $\phi(x)=\operatorname{ELU}(x)+1$.
 This kernel function ensures that the denominator remains positive to avoid the risk of division by zero, while the summation in the numerator also contains no negative terms.
-Inspired by techniques for stabilizing gradients in dot-product attention (Vaswani et al., 2017), we further incorporate scaling factors.
+Inspired by techniques for stabilizing gradients in dot-product attention [52], we further incorporate scaling factors.
 Specifically, the keys $\boldsymbol{K}$ and values $\boldsymbol{V}$ are scaled by $\sqrt{N}$ (where $N$ denotes the sequence length).
@@
 The final RELA is formulated as follows:
-(12) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)/\sqrt{N}\right)\left(\boldsymbol{V}_{n}/\sqrt{N}\right)\right)}{\phi(\boldsymbol{Q}_{m})\left(\frac{1}{N}\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}+\epsilon\right)}.$
+$\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}/\sqrt{N}\right)\right)}{\phi(\boldsymbol{Q}_{m})\frac{1}{N}\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}+\epsilon}.$ (12)
 In-Depth Analysis for Rotaty Position Encoding.
@@
 We first express it in its complex form:
-(13) $\boldsymbol{Q}_{m}=\left[\boldsymbol{Q}_{m}^{(1)},\boldsymbol{Q}_{m}^{(2)}\right]=\left[\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right],$
+$\boldsymbol{Q}_{m}=\left[\boldsymbol{Q}_{m}^{(1)},\boldsymbol{Q}_{m}^{(2)}\right]=\left[\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right],$ (13)
 where $\boldsymbol{Q}_{m}^{(1)}$ and $\boldsymbol{Q}_{m}^{(2)}$ indicate the components of $\boldsymbol{Q}_{m}$ expressed in the 2D coordinates.
 Based on Euler’s formula, $e^{im\theta}$ can be rewritten as follows:
-(14) $e^{im\theta}=\cos m\theta+i\sin m\theta.$
+$e^{im\theta}=\cos m\theta+i\sin m\theta.$ (14)
 Then, we have:
-(15) $\boldsymbol{Q}_{m}e^{im\theta}$ $=\left(\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right)(\cos(m\theta)+i\sin(m\theta))$
+$\boldsymbol{Q}_{m}e^{im\theta}$ $=\left(\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right)(\cos(m\theta)+i\sin(m\theta))$ (15)
 $=\left(\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)\right$. $\left.-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta)\right)+i\left(\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)\right).$
 Then, the result is re-expressed in the form of a real-valued vector:
-(16) $\boldsymbol{Q}_{m}e^{im\theta}$ $=[\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta),\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)]$
+$\boldsymbol{Q}_{m}e^{im\theta}=[$ $\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta),$ (16)
+$\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)]$
 $=\left(\begin{array}[]{cc}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{array}\right)\binom{\boldsymbol{Q}_{m}^{(1)}}{\boldsymbol{Q}_{m}^{(2)}}$.
 Evidently, the matrix $\left(\begin{array}[]{cc}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{array}\right)$ describes a rotational transformation applied to the embedded sequence.
-While the core rotational mechanism of RoPE effectively injects relative positional information, a critical design choice is the dimension upon which these rotations are applied.
-Conventionally, in multi-head attention architectures, RoPE is applied independently to each head’s query and key vectors, meaning the rotations operate on the head dimension $d$ (Su et al., 2024; Tian et al., 2024).
-This allows each head to potentially learn or leverage positional cues in its specialized subspace.
-Our choice deviates from this standard practice.
-We apply RoPE directly to the full embedding dimension $D$ of the queries and keys.
-The motivation is that imbuing the entire token representation with a unified positional signal at the embedding dimension $D$ level can lead to a more holistic and consistent understanding of sequence order throughout the model.
-Instead of each head independently re-interpreting or establishing positional context, all downstream computations operate on an input that is already position-aware across its full breadth.
-3.2.2. Local Shortcut and Gated Mechanism
+3.2.2 Local Shortcut and Gated Mechanism
 Attention mechanisms excel at long-range modeling, yet local behavior modeling is equally vital for long-term sequential recommendation systems, as user interactions often display strong local preferences (e.g., session-based behaviors) crucial for next-item prediction.
 However, linear attention variants falter here.
-Unlike Softmax attention’s sharp focus, linear attention’s smooth distribution tends to average features, obscuring crucial local cues (Han et al., 2023).
-Moreover, RoPE suffers from long-term decay (Su et al., 2024), meaning the attention score decreases as the relative distance increases.
+Unlike Softmax attention’s sharp focus, linear attention’s smooth distribution tends to average features, obscuring crucial local cues [13].
+Moreover, RoPE suffers from long-term decay [46], meaning the attention score decreases as the relative distance increases.
 Because of linear attention’s smoothing and RoPE’s distance decay, precisely modeling crucial local patterns for long-term sequential recommendation becomes difficult.
 To address these issues, we introduce a Local Shortcut to mitigate the long-term decay of RELA.
-Specifically, we use a causal 1D convolution (Fu et al., 2023) followed by a SiLU activation (Ramachandran et al., 2017) to capture local preferences in the user interaction sequence:
-(17) $\boldsymbol{S}=\operatorname{SiLU}\left(\text{CausalConv1d}\left(\boldsymbol{V}\right)\right),$
+Specifically, we use a causal 1D convolution [10] followed by a SiLU activation [41] to capture local preferences in the user interaction sequence:
+$\boldsymbol{S}=\operatorname{SiLU}\left(\text{CausalConv1d}\left(\boldsymbol{V}\right)\right),$ (17)
 where $\boldsymbol{V}$ denotes the input sequence, and $\boldsymbol{S}$ represents the output of the temporal convolution.
 Then $\boldsymbol{S}$ is connected with the output $\boldsymbol{O}$ of linear attention as a shortcut to form the final output $\boldsymbol{O}_{f}$ of RELA:
-(18) $\boldsymbol{O}_{f}=\boldsymbol{O}+\boldsymbol{S}.$
+$\boldsymbol{O}_{f}=\boldsymbol{O}+\boldsymbol{S}.$ (18)
 With the local shortcut, each query only focuses on several adjacent items instead of the whole sequence.
 This locality ensures that even if the linear attention values corresponding to two queries are the same, we can still get different outputs from different local preferences.
-The theoretical motivation behind such a design concerns the upper bound on the rank of the attention matrix in linear attention (Han et al., 2023).
+The theoretical motivation behind such a design concerns the upper bound on the rank of the attention matrix in linear attention [13].
 The rank of the attention matrix is bounded by the length of the sequence $N$ and the feature dimension $d$ for each head:
-(19) $\operatorname{rank}(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})^{\top})$ $\leq\min\{\operatorname{rank}(\phi(\boldsymbol{Q})),\operatorname{rank}(\phi(\boldsymbol{K}))\}$
+$\operatorname{rank}(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})^{\top})$ $\leq\min\{\operatorname{rank}(\phi(\boldsymbol{Q})),\operatorname{rank}(\phi(\boldsymbol{K}))\}$ (19)
 $\leq\min\{N,d\},$
@@
 18, the operation on the attention matrix can be reformulated as:
-(20) $\boldsymbol{O}_{f}$ $=\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\operatorname{CSC}(\boldsymbol{V})$
+$\boldsymbol{O}_{f}$ $=\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\operatorname{CSC}(\boldsymbol{V})$ (20)
 $=(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\boldsymbol{M}_{\operatorname{CSC}})\boldsymbol{V}=\boldsymbol{M}_{eq}\boldsymbol{V},$
@@
 To address the challenge of differentiating between temporary local preferences and genuine shifts in long-term user interests, we design a gated mechanism to intelligently modulate RELA’s output.
-We utilize a SiLU gate (Qin et al., 2023; Gu and Dao, 2024) and functioning as an output gate, learns to derive weights from the interaction sequence.
+We utilize a SiLU gate [39, 12] and functioning as an output gate, learns to derive weights from the interaction sequence.
 These weights then adjust the influence of different signals within the user behavior sequence processed by RELA.
@@
 The overall structure of the GRELA Block is shown as follows:
-(21) $\boldsymbol{H}_{g}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}\right),$
+$\boldsymbol{H}_{g}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}\right),$ (21)
 $\boldsymbol{H}_{1}$ $=\operatorname{SiLU}\left(\boldsymbol{W}_{1}\boldsymbol{H}_{g}+\boldsymbol{b}_{1}\right),$
@@
 where $\boldsymbol{H}_{g}$ is the input of the GRELA block, i.e., when $L=1$, the input $\boldsymbol{H}$ is from the embedding layer, $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{D\times D}$ are the weight matrices of the linear projection, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$ are the biases.
-Then a Hadamard product is performed on $\boldsymbol{H}_{1}$ and $\boldsymbol{H}_{2}$ followed by a linear projection and Droppath (Larsson et al., 2017) operation:
-(22) $\boldsymbol{H}_{3}=\operatorname{DropPath}\left(\boldsymbol{W}_{3}\left(\boldsymbol{H}_{1}\odot\boldsymbol{H}_{2}\right)+\boldsymbol{b}_{3}\right)+\boldsymbol{H},$
+Then a Hadamard product is performed on $\boldsymbol{H}_{1}$ and $\boldsymbol{H}_{2}$ followed by a linear projection and Droppath [22] operation:
+$\boldsymbol{H}_{3}=\operatorname{DropPath}\left(\boldsymbol{W}_{3}\left(\boldsymbol{H}_{1}\odot\boldsymbol{H}_{2}\right)+\boldsymbol{b}_{3}\right)+\boldsymbol{H},$ (22)
 where $\boldsymbol{W}_{3}\in\mathbb{R}^{D\times D}$ is the weight matrix, $\boldsymbol{b}_{3}$ is the bias.
-Meanwhile, due to the linear projection before the RELA, we can treat it as an input gate of linear attention like long-term memory network (LSTM), previous studies (Greff et al., 2017; Van Der Westhuizen and Lasenby, 2018) have demonstrated that the input gate plays a less significant role compared to the forget gate in LSTM.
+Meanwhile, due to the linear projection before the RELA, we can treat it as an input gate of linear attention like long-term memory network (LSTM), previous studies [11, 51] have demonstrated that the input gate plays a less significant role compared to the forget gate in LSTM.
 Inspired by this, we remove the linear projection for the value matrix of linear attention.
 To this end, we only use the weight matrices on query and key, reducing training costs.
-3.3. Multi-Layer Perceptron
-To capture the complex features, we further leverage a Multi-Layer Perceptron (MLP) (Kang and McAuley, 2018):
-(23) $\boldsymbol{H}_{4}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}_{3}\right),$
+3.3 Multi-Layer Perceptron
+To capture the complex features, we further leverage a Multi-Layer Perceptron (MLP) [19]:
+$\boldsymbol{H}_{4}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}_{3}\right),$ (23)
 $\boldsymbol{H}_{L}$ $=\operatorname{Drop}\left(\operatorname{Drop}\left(\operatorname{GELU}\left(\boldsymbol{H}_{4}\boldsymbol{W}_{4}+\boldsymbol{b}_{4}\right)\right)\boldsymbol{W}_{5}+\boldsymbol{b}_{5}\right)+\boldsymbol{H}_{3},$
@@
 $\boldsymbol{H}_{L}$ indicates the final output representation, and $\boldsymbol{H}_{4}$ denotes the output of layer normalization after the GRELA block.
-To stabilize the training process and enhance gradient flow through the network, we employ a residual connection (He et al., 2016).
+To stabilize the training process and enhance gradient flow through the network, we employ a residual connection [16].
 For generality, the subscript (0) indicates that the final user representation is obtained using a single GRELA layer.
-3.4. Prediction Layer and Training
+3.4 Prediction Layer and Training
 After the computation of $L$ GRELA blocks and the MLP, we get the final representation of behavior sequences with user interest.
@@
 This result is then scaled using the Softmax function to obtain the interaction probability score:
-(24) $\hat{\mathcal{Y}}=\operatorname{Softmax}(\boldsymbol{E}^{\top}\boldsymbol{H}_{L}).$
-Hence, following previous works (Kang and McAuley, 2018; Liu et al., 2024b), we expect to minimize the cross-entropy between the predicted recommendation results $\hat{\mathcal{Y}}$, and the ground truth items $\mathcal{Y}$:
-(25) $\mathcal{L}_{CE}\left(\mathcal{Y},\hat{\mathcal{Y}}\right)=\mathcal{Y}\log\left(\hat{\mathcal{Y}}\right)+\left(1-\mathcal{Y}\right)\left(1-\log\left(\hat{\mathcal{Y}}\right)\right).$
-3.5. Complexity Analysis
+$\hat{\mathcal{Y}}=\operatorname{Softmax}(\boldsymbol{E}^{\top}\boldsymbol{H}_{L}).$ (24)
+Hence, following previous works [19, 30], we expect to minimize the cross-entropy between the predicted recommendation results $\hat{\mathcal{Y}}$, and the ground truth items $\mathcal{Y}$:
+[…]
+3.5 Complexity Analysis
 We analyze the computational complexity of RELA (as shown in Eq.
@@
 Consequently, the overall time complexity of the RELA is $\mathcal{O}(Nd^{2})$.
-4. Experiments
-Table 3. Statistics of the experimental datasets.
+4 Experiments
+Table 1: Statistics of the experimental datasets.
 Datasets #Users #Items #Interactions Avg. UA Avg. IA Sparsity
@@
 ML-32M 200,949 43,885 31,921,467 158.9 727.4 99.63%
+LFM-1B 78,113 179,515 15,628,182 200.1 87.1 99.89%
 Netflix 120,378 16,087 21,767,054 180.8 1,353.1 98.86%
-4.1. Experimental Settings
-4.1.1. Datasets
-We conduct extensive experiments and verify the effectiveness of our approach on four real-world datasets, including MovieLens-1M, MovieLens-20M, MovieLens-32M  [^2] https://grouplens.org/datasets/movielens/ (dubbed as ML-1M, ML-20M, ML-32M) and Netflix  [^3] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data.
+4.1 Experimental Settings
+4.1.1 Datasets
+We conduct experiments and verify the effectiveness of our approach on five datasets, including MovieLens-1M, 20M, and 32M  [^1] https://grouplens.org/datasets/movielens/ (dubbed as ML-1M, ML-20M, ML-32M), LFM-1B  [^2] https://www.cp.jku.at/datasets/LFM-1b/, and Netflix  [^3] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data.
 The dataset statistics are shown in Table 1, where Avg.
@@
 Notably, for MovieLens datasets, we filter out users and items with fewer than 5 interactions.
+For LFM-1B datasets, we use the processed atomic files of LFM1B-Artists (merged) and filter out users with fewer than 50 and more than 500 interactions.
 Netflix is a huge-scale dataset, containing approximately 100 million interactions and five thousand Avg.
@@
 To simplify the training process, we filter out users with fewer than 50 interactions while retaining items with interactions ranging from 50 to 15,000 for this dataset.
-4.1.2. Baseline Models
-To verify the effectiveness and efficiency of RecGRELA, we make comparisons between RecGRELA and strong RNN-, Attention-, and Mamba-based baselines:
-  • LinRec (Liu et al., 2023): This model reduces the computational costs via linear attention in transformer-based models. We choose SASRec (Kang and McAuley, 2018) as the backbone of LinRec.
-  • LRURec (Yue et al., 2024): This method is based on linear recurrent units (Orvieto et al., 2023) for sequential recommendation.
-  • Mamba4Rec (Liu et al., 2024b): This is the first Mamba-based method, which explores applying the Mamba architecture in sequential recommendation tasks.
-  • EchoMamba4Rec (Wang et al., 2024): This method processes sequences both forward and backward to enhance the sequential representation with Mamba and frequency domain filter.
-  • EulerFormer (Tian et al., 2024): This method proposes a transformer variant with complex vectors based on Euler’s formula and a contrastive learning method to improve contextual representations.
-  • RecBLR (Liu et al., 2024a): This model shows a behavior-dependent linear recurrent unit to enhance user behavior modeling with a hardware-aware parallel scan approach.
-  • SIGMA (Liu et al., 2025): This model employs a flipped Mamba to build a bidirectional framework to improve short sequence representations and contextual modeling.
-Table 4.
-Performance comparisons on ML-1M, ML-20M, ML-32M, and Netflix datasets.
+4.1.2 Baseline Models
+To verify the effectiveness and efficiency of RecGRELA, we make comparisons between RecGRELA and strong RNN-, Attention-, and Mamba-based baselines: LinRec [31]: This model reduces the computational costs via linear attention in transformer-based models.
+We choose SASRec [19] as the backbone of LinRec.LRURec [62]: This method is based on linear recurrent units [36] for sequential recommendation.
+Mamba4Rec [30]: This is the first Mamba-based method, which explores applying the Mamba architecture in sequential recommendation tasks.
+EchoMamba4Rec [56]: This method processes sequences both forward and backward to enhance the sequential representation with Mamba and frequency domain filter.
+EulerFormer [50]: This method proposes a transformer variant with complex vectors based on Euler’s formula and a contrastive learning method to improve contextual representations.
+RecBLR [29]: This model shows a behavior-dependent linear recurrent unit to enhance user behavior modeling with a hardware-aware parallel scan approach.
+SIGMA [32]: This model employs a flipped Mamba to build a bidirectional framework to improve short sequence representations and contextual modeling.
+Table 2: Performance comparisons on ML-1M, ML-20M, ML-32M, and Netflix datasets.
 Best and second-best results per row are shown in bold and underlined format, respectively.
-”Improv.” denotes the percentage improvement of the best-performing model over the second-best for each metric.
+"Improv." denotes the percentage improvement of the best-performing model over the second-best for each metric.
+The $\star$” denotes the statistical significance ($p$ < 0.05) of the results of MegaRec compared to the best baseline.
 Dataset Metric LinRec LRURec Mamba4Rec EchoMamba EulerFormer RecBLR SIGMA RecGRELA Improv.
@@
 MRR@10 0.1090 0.1015 0.1062 0.1083 0.1003 0.1073 0.1116 0.1175 5.29%
+LFM-1B HR@5 0.0755 0.0732 0.0768 0.0781 0.0745 0.0778 0.0834 0.0860 3.12%
+NDCG@5 0.0421 0.0401 0.0435 0.0448 0.0410 0.0447 0.0512 0.0531 3.71%
+MRR@5 0.0315 0.0298 0.0322 0.0339 0.0305 0.0339 0.0407 0.0424 4.18%
+HR@10 0.1345 0.1288 0.1352 0.1380 0.1333 0.1420 0.1421 0.1460 2.74%
+NDCG@10 0.0632 0.0603 0.0640 0.0655 0.0621 0.0702 0.0700 0.0724 3.13%
+MRR@10 0.0465 0.0441 0.0473 0.0481 0.0456 0.0482 0.0484 0.0502 3.72%
 Netflix HR@5 0.0900 0.0712 0.0809 0.0827 0.0781 0.0873 0.0897 0.0922 2.44%
@@
 MRR@10 0.0582 0.0484 0.0532 0.0553 0.0495 0.0567 0.0582 0.0611 4.98%
-Figure 3. RecGRELA vs. baselines in terms of the GPU memory and FLOPs of the training stage on the ML-1M dataset.
-4.1.3. Evaluation Protocols
+Figure 3: RecGRELA vs. baselines in terms of the GPU memory and FLOPs of the training stage on the ML-1M dataset.
+4.1.3 Evaluation Protocols
 We evaluate performance using three standard metrics: Hit Rate (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR).
 These metrics are assessed based on top-$K$ recommendation lists, specifically using cutoff values $K$ = 5 and $K$ = 10, denoted as HR@5, NDCG@5, MRR@5, HR@10, NDCG@10, and MRR@10, respectively.
-4.1.4. Implementation Details
-We utilize the RecBole library  [^4] https://github.com/RUCAIBox/RecBole (Zhao et al., 2021; Xu et al., 2023), a unified framework for implementing sequential recommendations.
+4.1.4 Implementation Details
+We utilize the RecBole library  [^4] https://github.com/RUCAIBox/RecBole [66, 59], a unified framework for implementing sequential recommendations.
 To ensure a fair comparison, we fix the embedding size to 64, the hidden size of the MLP to 256, and the training batch size to 2048 for all the models.
@@
 We set the Dropout and Droppath rate to 0.1.
-We use Adam (Kingma, 2014) with a learning rate of 0.001 for optimizing our model.
+We use Adam [21] with a learning rate of 0.001 for optimizing our model.
 The maximum sequence length $N$ is set to 200 for all the datasets.
@@
 All the experiments are conducted on a single NVIDIA RTX4090.
-4.2. Overall Performance
-Table 4 reports the recommendation performance of RecGRELA compared to seven baselines on four datasets. From the table, we find several observations:
+4.2 Overall Performance
+Table 2 reports the recommendation performance of RecGRELA compared to seven baselines on five datasets. From the table, we find several observations:
   • Our RecGRELA outperforms all competing transformer-based, RNN-based, and Mamba-based baselines, with improvements ranging from 0.61% to 4.98%.
@@
 This can also be attributed to introducing RoPE, which enhances the model’s ability to capture positional information.
-4.3. Efficiency Analysis
+4.3 Efficiency Analysis
 We compare the training efficiency of our RecGRELA with state-of-the-art sequential recommendation models in terms of GPU memory and FLOPs.
@@
 […]
-Table 5. Sensitivity analysis on different numbers of layers. The sequence length is set to 200. Time indicates the training time per epoch, and Memory indicates the GPU memory.
+Table 3: Sensitivity analysis on different numbers of layers. The sequence length is set to 200. Time indicates the training time per epoch, and Memory indicates the GPU memory.
 #Layers HR@5 NDCG@5 MRR@5 HR@10 NDCG@10 MRR@10 Time Memory
@@
 6 0.1776 0.1109 0.0948 0.3242 0.1912 0.1506 113.43s 17.68GB
-Figure 4. The effectiveness of each variant of RecGRELA on ML-1M.
-4.4. Ablation Study
-4.4.1. Core Components
+Figure 4: The effectiveness of each variant of RecGRELA on ML-1M.
+4.4 Ablation Study
+4.4.1 Core Components
 To investigate how different components contribute to our design, we evaluate RecGRELA variants by removing the key components one by one and implementing some modifications as follows:
@@
 This aligns with the observations in Section 4.2.
-Figure 5. The performance of different position encodings and activation functions.
-4.4.2. Position Encoding and Activation Function
+Figure 5: The performance of different position encodings and activation functions.
+4.4.2 Position Encoding and Activation Function
 We further investigate the impact of position encoding strategies on RecGRELA.
@@
 5, the SiLU activation function outperforms ReLU and GELU across all metrics, demonstrating the superior performance of SiLU in capturing complex user behavior and enhancing the model’s representational capacity.
-Figure 6. Impact of different stacked layers on computational costs. We set the sequence length to 100 to avoid GPU memory errors.
-4.5. Sensitivity Analysis
+Figure 6: Impact of different stacked layers on computational costs. We set the sequence length to 100 to avoid GPU memory errors.
+4.5 Sensitivity Analysis
 In this section, we conduct experiments on the ML-1M dataset to analyze the influence of three hyperparameters: the sequence length $N$, the number of stacked GRELA blocks $L$, the number of attention heads $h$, and the Dropout rates.
-4.5.1. The Number of Layers
+4.5.1 The Number of Layers
 The results in Table 3 reveal that stacking more RecGRELA layers fails to consistently enhance recommendation quality while markedly degrading computational efficiency, primarily due to the overfitting problem of multiple RecGRELA layers.
@@
 6, the growth trend of computational occupation of RecGRELA is relatively gradual, whereas that of SIGMA shows a steeper increase, indicating that RecGRELA exhibits better scaling capability when the training requires more layers.
-Figure 7.
-Impact of sequence length and attention head variations on computational costs.
+Figure 7: Impact of sequence length and attention head variations on computational costs.
 When investigating the impact of sequence length, we set the attention head to 2 for EulerFormer and 4 for RecGRELA.
@@
 Other hyperparameters are the same as the main experiments.
-4.5.2. The Number of Sequence Length
+4.5.2 The Number of Sequence Length
 To examine training cost variations across sequence lengths in long-term SRSs, we evaluate maximum sequence lengths from 100 to 200 with 25-step increments.
@@
 This suggests that while the model performs consistently well with shorter sequences, capturing longer-range dependencies present in sequences of length 200 yields a clear benefit in terms of recommendation performance.
-4.5.3. The Number of Attention Heads
+4.5.3 The Number of Attention Heads
 We evaluate the impact of the number of attention heads on model efficiency.
@@
 Therefore, four attention heads represent the optimal setting identified in our experiments.
-4.5.4. The Dropout and Droppath Rates
+4.5.4 The Dropout and Droppath Rates
 Regularization techniques like Dropout and Droppath prevent overfitting and improve model generalization.
@@
 This highlights the importance of carefully tuning regularization hyperparameters, as overly aggressive regularization can harm performance more than it helps prevent overfitting in sequential recommendation.
-Figure 8. Influence of the sequence length, head numbers, Dropout, and Droppath on various metrics.
-4.6. Visualization Analysis
+Figure 8: Influence of the sequence length, head numbers, Dropout, and Droppath on various metrics.
+Figure 9: Heatmap of output feature maps from the last layer.
+4.6 Visualization Analysis
 To further investigate the impact of the proposed Local Shortcut on the representational capacity of our model, particularly regarding the theoretical motivation to increase feature matrix rank, we conduct a visualization analysis on the ML-1M dataset.
@@
 It helps the model capture immediate local preferences and crucial long-term dependencies within the user’s interaction sequence.
-Figure 9. Heatmap of output feature maps from the last layer.
-5. Related Work
-5.1. Sequential Recommendation
-Earlier fundamental models of sequential recommendation focus on using Convolutional Neural Networks (Tang and Wang, 2018; Yuan et al., 2020) and Recurrent Neural Networks (RNNs) (Hidasi, 2016; Li et al., 2017).
-Due to RNNs’ performance limits and poor parallel processing, recent sequential recommendation studies have increasingly turned to Transformer-based architectures (Wu et al., 2020; Li et al., 2021; Du et al., 2023; Fan et al., 2022; Zhang et al., 2023; Zhao et al., 2023; Li et al., 2023a; Shi et al., 2024) or MLPs (Long et al., 2024; Zhou et al., 2022; Li et al., 2023b, 2022; Gao et al., 2024; Jiang et al., 2024) for better efficiency and scalability.
-The most representative work is SASRec (Kang and McAuley, 2018), which is the first to use the self-attention mechanism for capturing user interest features from interaction sequences.
-BERT4Rec (Sun et al., 2019) adopts a deep encoder-based bidirectional model with self-attention to model user behavior sequences, achieving strong recommendation performance.
+5 Related Work
+5.1 Sequential Recommendation
+Earlier fundamental models of sequential recommendation focus on using Convolutional Neural Networks [49, 61] and Recurrent Neural Networks (RNNs) [18, 25].
+Due to RNNs’ performance limits and poor parallel processing, recent sequential recommendation studies have increasingly turned to Transformer-based architectures [57, 28, 5, 8, 64, 65, 24] or MLPs [33, 67, 27] for better efficiency and scalability.
+The most representative work is SASRec [19], which is the first to use the self-attention mechanism for capturing user interest features from interaction sequences.
+BERT4Rec [47] adopts a deep encoder-based bidirectional model with self-attention to model user behavior sequences, achieving strong recommendation performance.
 However, the Transformer architecture suffers from quadratic computational complexity with respect to sequence length due to the dot-product attention operation, which limits its efficiency for sequence modeling tasks.
-Recently, many models with linear computational complexity have emerged as alternatives to the original Transformer architecture, such as linear attention (Liu et al., 2023; Tian et al., 2024), linear RNNs (Liu et al., 2024a; Yue et al., 2024), and state space models (Liu et al., 2024b; Qu et al., 2024; Fan et al., 2024).
-While LinRec (Liu et al., 2023) adopts linear attention to reduce the computational complexity of the dot-product attention, they still rely on learnable position encodings to capture positional information.
-EulerFormer (Tian et al., 2024) introduces a novel Euler transformation with a rotary form for unified modeling of both semantic and positional difference, yet uses the dot-product attention, which results in computational costs comparable to LinRec.
+Recently, many models with linear computational complexity have emerged as alternatives to the original Transformer architecture, such as linear attention [31, 50], linear RNNs [29, 62], and state space models [30, 40, 6].
+While LinRec [31] adopts linear attention to reduce the computational complexity of the dot-product attention, they still rely on learnable position encodings to capture positional information.
+EulerFormer [50] introduces a novel Euler transformation with a rotary form for unified modeling of both semantic and positional difference, yet uses the dot-product attention, which results in computational costs comparable to LinRec.
 To overcome these specific limitations, RecGRELA combines linear attention with rotary position encodings and introduces a gated mechanism to model both local and long-term user interests.
-5.2. Linear Attention
-As opposed to dot-product attention, linear attention (Katharopoulos et al., 2020) replaces the Softmax with separate kernel functions, thereby reducing computational complexity to $\mathcal{O}(N)$ through a change in computation order.
+5.2 Linear Attention
+As opposed to dot-product attention, linear attention [20] replaces the Softmax with separate kernel functions, thereby reducing computational complexity to $\mathcal{O}(N)$ through a change in computation order.
 Though efficient, designing a linear attention module as effective as softmax attention is a nontrivial problem.
-Although linear attention has achieved some success in other domains, such as computer vision (Shen et al., 2021; Qin et al., 2022; Han et al., 2024a, b) and neural language processing (Wang et al., 2020; Qin et al., 2023; Yang et al., 2024).
-In sequential recommendation, despite existing linear attention-based methods (Fan et al., 2021; Liu et al., 2023) decreasing the computational complexity, they cannot effectively model position encoding within behavior sequence and fail to capture local preference in long-term interactions.
+Although linear attention has achieved some success in other domains, such as computer vision [44, 38, 14, 15] and neural language processing [55, 39, 60].
+In sequential recommendation, despite existing linear attention-based methods [7, 31] decreasing the computational complexity, they cannot effectively model position encoding within behavior sequence and fail to capture local preference in long-term interactions.
 This motivates us to design a more efficient and effective linear attention-based module to address these issues.
 In response, we present RecGRELA, which aims to realize the full potential of linear attention in sequential recommendation by overcoming these critical challenges in positional awareness and local preference modeling.
-6. Conclusion and Future Work
+6 Conclusion and Future Work
 In this work, we propose RecGRELA, a new paradigm for long-term sequential recommendation.
@@
 To address the difficulty of the model in perceiving and adapting to dynamic user preference shifts, we propose a SiLU-based gating mechanism that dynamically balances local and long-term representations.
-Extensive experiments on four benchmark datasets show that RecGRELA outperforms state-of-the-art methods based on RNNs, Transformers, and Mamba architectures, achieving superior performance with significantly lower memory overhead.
+Extensive experiments on five benchmark datasets show that RecGRELA outperforms state-of-the-art methods based on RNNs, Transformers, and Mamba architectures, achieving superior performance with significantly lower memory overhead.
 In the future, we plan to explore the application of RecGRELA in other recommendation scenarios, such as session-based and multi-modal recommendations.
 Since linear attention shows great potential in various domains, we also plan to explore more efficient attention mechanisms for sequential recommendation tasks.
-Acknowledgements.
-This work is funded by the National Natural Science Foundation of China (Grant No.
-72074036, 62072060) and supported by the Natural Science Foundation of Chongqing, China, No.
-CSTB2024NSCQ-MSX0701, CSTB2024NSCQ-MSX0617.
+[…]